### v1

# A Comprehensive Benchmark for RNA 3D Structure-Function Modeling

Abstract
The RNA structure-function relationship has recently garnered significant attention within the deep learning community, promising to grow in importance as nucleic acid structure models advance. However, the absence of standardized and accessible benchmarks for deep learning on RNA 3D structures has impeded the development of models for RNA functional characteristics.
In this work, we introduce a set of seven benchmarking datasets for RNA structure-function prediction, designed to address this gap. Our library builds on the established Python library rnaglib, and offers easy data distribution and encoding, splitters and evaluation methods, providing a convenient all-in-one framework for comparing models. Datasets are implemented in a fully modular and reproducible manner, facilitating for community contributions and customization. Finally, we provide initial baseline results for all tasks using a graph neural network.
RNA Structure-Function Relationship, Graph Learning, Benchmarks, RNA 3D Structure, Computational Biology, RNA Structural Modeling
Source code: github.com/cgoliver/rnaglib
Documentation: rnaglib.org
1 Introduction
Recent years have witnessed the advent of deep learning methods for structural biology culminating in the award of the Nobel Prize in Chemistry. AlphaFold (Jumper et al., 2021) revolutionized protein structure prediction, equipping the field with millions of new structures. Breakthroughs go beyond structure prediction, notably in protein design (Watson et al., 2023; Dauparas et al., 2022), drug discovery (Schneuing et al., 2024; Corso et al., 2022) or fundamental biology (van Kempen et al., 2022). While it is tempting to attribute the success of these methods to the increase in available structural data caused by AlphaFold, most of the methods are actually not reliant on them. Instead, it seems that these breakthroughs result from progress in training neural encoders that directly model protein structures (Jing et al., 2020; Zhang et al., 2022b; Gainza et al., 2020; Wang et al., 2022). This progress is in turn rooted in solid competitions (CASP, CAPRI), and benchmarks (Townshend et al., 2021a; Kucera et al., 2023; Zhu et al., 2022; Jamasb et al., 2024; Notin et al., 2023). By setting clear goals, such benchmarks are the foundation for the development of structure encoders. Yet to date, structure-function benchmarks have focused on proteins.
Ribonucleic acids (RNAs) are a large family of molecules which support biological functions along every branch of the tree of life. Besides messenger RNAs, non-coding RNAs carry out biological functions by adopting complex 3D folds (Cech & Steitz, 2014) like proteins do and take up diverse roles in cellular functions, including gene regulation, RNA processing, and protein synthesis (Statello et al., 2021). However, our understanding of non-coding RNAs and their functions remains limited. This can be largely attributed to the negatively charged nature of RNA backbones, which makes it flexible and limits the availability of high-resolution RNA structures, and imposes significant modeling challenges. Another predominant challenge to a functional understanding of RNA 3D structure lies in the lack of infrastructure for the development and evaluation of function prediction models. In this work, we propose a benchmarking suite to act as this facilitating framework.
Our key contributions include:
  • Seven tasks related to RNA 3D structure that represent various biological challenges. Each task consists of a dataset, a splitting strategy, and an evaluation method, laying the ground for comparable, reproducible model development.
  • End-to-end reproducible and modular access to task data. Modular annotators, filters and splitting strategies, both novel and from existing literature, facilitate the addition of new tasks by other researchers across fields.
2 Related Work
2.1 Protein Focused Benchmarking
Classic tasks on proteins appeared independently in unrelated papers, such as GO term and EC number prediction (Gligorijević et al., 2021), fold classification (Hou et al., 2018), binding site detection and classification (Gainza et al., 2020) or binding affinity regression (Wang et al., 2005a). To our knowledge, ATOM3D (Townshend et al., 2021a) was the first systematic benchmark for molecular systems, albeit heavily focused on proteins. Similar, more comprehensive tools were then proposed, such as ProteinShake (Kucera et al., 2023) ProteinWorkshop (Jamasb et al., 2024) and TorchDrug (Zhu et al., 2022), that unify the above tasks and lower the barrier to develop protein structure encoders. ProteinGym (Notin et al., 2023) addresses the evaluation of protein mutation effects, while Buttenschoen et al. (2024); Kovtun et al. (2024); Durairaj et al. (2024) address protein interactions. These works contain notable efforts to scale datasets using predicted structures and to propose biologically appropriate splitting strategies.
2.2 RNA Structural Datasets and Benchmarking
In the realm of RNA 3D structure based modeling infrastructure, three papers propose cleaned datasets with the objective of facilitating machine learning. RNANet (Becquey et al., 2021) proposes a dataset joining RNA structures with their corresponding sequence alignment. RNAsolo (Adamczyk et al., 2022) provides access to cleaned RNA files in various formats through a web interface. Finally, RNA3DB (Szikszai et al., 2024) offers a curated dataset specifically designed for RNA structure prediction models. None of these methods propose benchmark tasks to compare RNA modeling and learning methods. Moreover, some datasets along with splits were independently proposed in several work, pertaining to small-molecules binding sites (Wang et al., 2018), or virtual screening (Panei et al., 2022; Carvajal-Patiño et al., 2025). The inverse folding field has seen a particular breath of advances in (Joshi et al., 2024; Tan et al., 2023, 2025; Huang et al., 2024; Wong et al., 2024). Recent benchmarking suites like BEACON (Ren et al., 2024) provide extensive tools for models on RNA, but do not rely on structural data or focus on non-coding RNAs.
2.3 Structure-Based RNA Models
While most deep learning models on RNA focus on secondary structure prediction and sequence-level tasks, some structure-based models have been developed. RBind (Wang et al., 2018) proposed learning on RNA structures using residue graph representations, followed by others integrating sequence features in the learning (Su et al., 2021; Wang et al., 2023). RNAmigos and RNAmigos2 (Oliver et al., 2020; Carvajal-Patiño et al., 2025) proposed incorporating non-canonical interactions in the graph, in conjunction with metric learning pretraining. In a follow-up work (Wang et al., 2023), a similar method for binding site prediction was used. Finally, gRNAde (Joshi et al., 2024) adapted the popular GVP (Jing et al., 2021) protein encoder, which uses the atomic position in the message passing algorithm, to RNA.
Despite manifold efforts at deep learning-based modeling RNA 3D structure, we lack the means to systematically compare models and onboard new practitioners.
3 Tools to assemble tasks
This section describes the full process of building a task: data collection, quality filtering, cleanup, and splitting.
Table 1: Description of the proposed datasets and biologically-relevant tasks.
Task Name Short description Novel Expanded
1. RNA-GO Classify RNAs into one or more frequent function. $\surd$
2. RNA-IF Find a sequence that folds into this structure. $\surd$
3. RNA-CM Predict which RNA residues are chemically modified from its geometry. $\surd$
4. RNA-Prot Predict which RNA residues are interacting with a protein. $\surd$
5. RNA-Site Predict which RNA residues are interacting with a small-molecule protein. $\surd$
6. RNA-Ligand Classify an RNA binding site, based on the ligand it binds to. $\surd$
7. RNA-VS Predict the affinity of a small molecule for an RNA binding site, and use it for Virtual Screening (VS). $\surd$
Data collection and annotation
Our data originates from the RCSB-PDB DataBank (Kouranov et al., 2006), where we fetch all 3D structures containing RNA. In some tasks, filter down the set of RNA-containing PDB structures with the subset proposed in (Leontis & Zirbel, 2012) and later referred to as bgsu. We annotate each system with RNA-level features such as its resolution, and residue-level features such as their coordinates, the presence of interacting compounds, and the amount of protein atoms in the residues’ vicinities.
Structure partitioning and quality filters
RNA molecules present in PDB files exhibit a bimodal distribution over the number of residues (see figures 1(a) and 1(b)). Many systems have less than 300 residues, while a few others (mostly ribosomal structures) have several thousands. To avoid discarding entire systems, we split whole systems into sub-structures by sets of connected components in the graph formed by all base pairings and backbone connections. Concretely, this allows independent RNA fragments appearing in the same PDB file to be treated as multiple systems; the resulting decrease in system size leads to a reduction in required computing power for downstream steps, such as similarity-based splitting.
We implement resolution and size filters, with default cutoff values including systems below 4Å resolution and with a length of 15 to 500 residues. In tasks, the upper size cutoff is used to limit computational expenses. We also provide a protein content filter which removes RNA structures that are heavily structured through an interaction with a protein. This crucial filter has been overlooked in most of existing RNA structural datasets. In addition, we implement the drug-like filter introduced in Hariboss (Panei et al., 2022) for small molecules binding to RNA.
Minimizing data leaks
Because biological data points are often related in terms of evolutionary history and topological characteristics, insufficient rigor in splitting can lead to severe data leaks. Unsurprisingly, numerous examples of data leakage have pestered structural biology learning methods, one famous example being the first PDBbind dataset that was shown to incorporate several severe biases (Wang et al., 2005b; Volkov et al., 2022). Efforts like the PINDER database (Kovtun et al., 2024) have set a standard in the field for model generalisation assessment.
In our work, we implement the computation of sequence-based or structure-based similarity matrices, with CD-HIT (Fu et al., 2012) and US-align (Zhang et al., 2022a) respectively. Given a similarity matrix $S$, computed over a set of RNA molecules $\mathcal{R}$, and a threshold $\theta$, we introduce the matrix $S^{\theta}$ defined as $S^{\theta}_{i,j}=\delta\{S_{i,j}\geq\theta\}$, where $\delta$ is the indicator function. We propose a clustering algorithm that considers $(\mathcal{R},S^{\theta})$ as a graph, and returns connected components as clusters. This ensures that any pair of points in different clusters has a maximal similarity of $\theta$.
Redundancy filtering and dataset splitting
Starting with clusters, we propose a redundancy removal algorithm that selects the element with the highest resolution for each cluster. In most tasks, we first apply this algorithm at a sequence similarity threshold of 0.9. This stringent threshold discards copies of systems with minimal variations that are common when getting the structure of a given system in slightly different conditions. Then, we apply another structure-based redundancy threshold of 0.8.
In addition, we propose a novel splitting algorithm that starts from clustered data, but uses a less conservative structural similarity threshold (0.5 unless mentioned otherwise). Then, we aggregate those clusters together to form splits of a certain size, optionally following a label balancing secondary objective, using the linear programming PuLP (Mitchell et al., 2011). This prevents data leakage between splits, while allowing split stratification based on labels.
Metrics computation
Given the type of prediction target (e.g. binary classification, multi-label, etc.) , each task implements a standardized way to compute evaluation metrics. These are specified for each task in Section 4.
Anatomy of a task
A “task” is now simply the collection of the aformentioned components with the choices relevant to a particular biological problem. That is, each task consists of (i) a dataset drawn from our 3D structure database and processed using our annotations, partitions and filters, with redundancy removing (ii) fixed splits based on sequence, structure or existing literature, and (iii) a well-defined evaluation protocol. This modular design will help other practitioners propose additional benchmark tasks, and lower the barrier to entry for training models on RNA structure.
4 Tasks on RNA 3D Structure
Our tasks suite explores various dimensions of RNA structural research briefly summarized in Table 1. We offer seven tasks among which three are based on previous research and offer publicly available split datasets (RNA-IF (Joshi et al., 2024), RNA-Site (Su et al., 2021) and RNA-VS (Carvajal-Patiño et al., 2025)). Moreover, for two of these three tasks, we propose an enhanced, expanded version, resulting in a total of nine separate datasets.
Armed with an RNA structure, the main focus of each of the subsequent tasks is to either predict aspects of the RNA’s biological function, or to reverse the direction and predict the sequence from which a given structure arises (RNA design). We provide an overview of these tasks from a machine learning point of view relating to the type of task and dataset size in Table 2.
Table 2: Description of the ML features of each of our tasks. The number in parenthesis refers to the number of class, in the multi-class setting. We include a comparison of the datasets we introduce with existing ones.
Task name Level Type Dataset split sizes
1. RNA-GO RNA Multi-class (5) 349-75-75
2. RNA-IF Residue Multi-class (4) 1700-448-581
   gRNAde (Joshi et al., 2024) 11183-528-235
3. RNA-CM Residue Binary 138-29-30
4. RNA-Prot Residue Binary 881-189-189
6. RNA-Site Residue Binary 157-34-33
   RNASite (Su et al., 2021) 53-6-17
6. RNA-Ligand Pocket Multi-class (3) 203-43-44
7. RNA-VS (Carvajal-Patiño et al., 2025) Pocket Regression 304-34-65
4.1 RNA-GO: Function Tagging
(Definition): This is a multi-label classification task where input RNA is mapped to a sequence of 5 possible labels representing molecular functions.
(Context): Effective function prediction models have the capacity to uncover new structure-function connections. The Gene Ontology (GO) (Ashburner et al., 2000) was developed to associate a function to a gene, and thus, indirectly to the RNA or protein it encodes. It resulted in discrete functional categories called GO terms. Predicting the GO term from a protein structure was proposed as a task in (Gligorijević et al., 2021) and has since been regularly used as a benchmark task. A mapping between RNA sequences and GO terms is available in Rfam (Griffiths-Jones et al., 2003; Ontiveros-Palacios et al., 2025), a database of non-coding RNA families which are manually annotated with GO terms.
(Processing): We extracted all GO terms corresponding to an RNA structure. Then, we filtered these terms based on their frequency, removing the two most frequent labels (ribosomal systems and tRNAs), as well as infrequent ones (less than 50 occurrences). We then grouped together GO-terms that were fully correlated, which resulted in five classes. Finally, we discarded RNA fragments with less than 15 residues, resulting in 501 systems. We split these systems based on a sequence similarity cutoff, following Gligorijević et al. (2021). Accuracy, AuROC and the Jaccard index are the metrics computed for evaluation.
4.2 RNA-IF: Molecular Design
(Definition): This is a residue-level, classification task where given an RNA structure, we mask all residue identities (by keeping only the coordinates of the backbone) and aim to learn the mapping from structure to sequence.
(Context): One avenue for designing new macromolecules is to design a coarse grained tertiary structure first, for instance scaffolding a structural motif, and then design a sequence that will fold into this given structure. This second step is denoted as Inverse Folding (IF), since it maps a known structure to an unknown sequence. Inverse folding is a well-established task in protein research with many recent breakthroughs (Watson et al., 2023). It has recently been adapted for RNA with classical models such as Rosetta (Leman et al., 2020a). Learning based approaches were pioneered by gRNAde (Leman et al., 2020a; Joshi et al., 2024) and other papers specifically address backbone generation (Anand et al., 2024) and protein-binding RNA design (Nori & Jin, 2024).
(Processing): We gather all connected components in our 3D structure database with a size between 15 and 300. Then, we removed identical sequences using a redundancy removal step with CD-HIT with a cutoff of 0.99. We cluster the data on structural similarity with a threshold of 0.9, and split these while ensuring a maximal similarity of 0.5. We also provide datasets and splits from gRNAde (Joshi et al., 2024), that were obtained in a similar manner. Our dataset differs by allowing multi-chain systems, a stricter size cutoff, and a duplicates filter, leading to a reduced dataset size. This task is evaluated with sequence recovery.
4.3 RNA-CM: Dynamic function modulation.
(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict which, if any, of its residues are chemically modified.
(Context): In addition to the four canonical nucleotides, more than 170 modified nucleotides were found to be integrated in RNA polymers (Boccaletto et al., 2022). Multiple functions of a diverse range of ncRNA have been shown to be directly dependent on such chemical modifications (Roundtree et al., 2017). We propose to detect such chemical modifications from subtle perturbations of the canonical geometry of the RNA backbones. Training a predictor on RNA structures is not the most straightforward way of detecting chemical modifications, but may uncover function-modulating modifications that hitherto went unnoticed.
(Processing): To build this task, we start from the whole dataset, partition it into connected components and apply our size filter. Then, we filter for systems that include modified residues, relying on PDB annotations that flag such modified residues. We apply our default redundancy removal and splitting strategies, which results in 185 data points. Performance is evaluated with accuracy and AuROC.
4.4 RNA-Prot: Biological complex modeling.
(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether a protein residue is closer than 8Å to any of its residues.
(Context): RNAs and proteins often bind to form a functional complex. Such complexes are involved in crucial cell processes, such as post-transcriptional control of RNAs (Glisovic et al., 2008). RNA structure is often, but not always, heavily disrupted upon interaction with a protein. We expect this task to be an easier version of the RNA-CM task.
(Processing): To build this task, we start from the bgsu non-redundant dataset, partition it into connected components and apply our size and resolution filters. We remove systems originating from ribosomes that are fully encapsulated in a complex. Then, we retain only RNA interacting with a protein and apply our default redundancy removal and splitting strategies, resulting in 1251 data points. Performance is evaluated with accuracy and AuROC.
Tasks for RNA drug design
Further tasks aim to address RNA small-molecule drug-design. RNA is increasingly recognized as a promising family of targets for the development of novel small molecule therapeutics  (Falese et al., 2021; Haga & Phinney, 2023; Disney, 2019; Abulwerdi et al., 2019). Being able to target RNA would drastically increase the size of the druggable space, and propose an alternative for overused protein targets in scenarios where they are insufficient. In addition, they represent a therapeutic avenue in pathologies where protein targets are absent, such as in triple-negative breast cancer (Xu et al., 2020).
4.5 RNA-Site: Drug target detection.
(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether a ligand is closer than 6Å to any of its residues.
(Context): The classical flow of a structure-based drug discovery pipeline starts with the identification of relevant binding sites, which are parts of the structure likely to interact with ligands, or of particular interest for a phenotypical effect. The structure of the binding site can then be used to condition the quest for small molecule binders, for instance using molecular docking (Ruiz-Carmona et al., 2014). The framing of this problem as a machine learning task for RNA was introduced in Rbind (Wang et al., 2018).
(Processing): We provide datasets and predefined splits from RNASite (Su et al., 2021), which are also utilized by other tools in the field such as RLBind (Wang et al., 2023). This dataset contains 76 systems obtained after applying stringent clustering on an older version of the PDB.
In addition to providing the existing dataset existing in the literature, we propose a larger, up-to-date dataset. We start by following similar steps as for RNA-Prot (without removal of ribosomal systems). We then include a drug-like filter on the small-molecule side. Finally, we include a protein-content filter, removing systems with more than 10 protein atom neighbors, ensuring that the binding is modulated by RNA only. We use the default redundancy removal and splitting, resulting in 458 systems. The predictions are evaluated using accuracy and AuROC.
4.6 RNA-Ligand: Pocket categorization
(Definition): This is a binding site-level, multi-class classification task where the structure of an RNA binding site is classified according to the partner it accommodates.
(Context): Equipped with a binding site, one wants to use its structure to characterize its potential binders. On proteins, the Masif-Ligand tasks (Gainza et al., 2020) gathers all binding sites bound to the seven most frequent co-factors, and aims to classify them based on their partner. Inspired by this work we propose the RNA-Ligand task. To keep sufficiently many examples per task, we only retained the three most frequent classes. This task is less ambitious than training a molecular docking surrogate and can help understanding the potential modulators of a given binding site.
(Processing): Starting with similar steps as the RNA-Site, we obtain a set of structures that display RNAs in interaction with one or more drug-like small molecules. We then proceed to extracting the context of the binding pocket by seeding two rounds of breadth-first search with all residues closer than 8Å to an atom of the binder. This results in all binding pockets in our database and we group these pockets with a sequence clustering.
Then, our aim is to find the most frequent ligands binding in non-redundant pockets. To that end, we gather a dataset of binding site clusters that bind to one ligand only. We then compute the most frequent ligands among this dataset and retain only the top three. Subsequently, we discard all binding events to other small molecules, and add clusters that bind only one of the top three ligands to our dataset. This dataset is split on structural similarity and evaluated on MCC and AuROC.
4.7 RNA-VS: Drug screening
(Definition): This is a binding site-level, regression task. Given the structure of an RNA binding site and a small molecule represented as a molecular graph, the goal is to predict the affinity of their binding.
(Context): Beyond classification into fixed categories, virtual screening aims to score compounds based on their affinity to a binding pocket. This task is ubiquitous in drug design as it helps selecting the most promising compounds to be further assayed in the wet-lab. Our last task implements a virtual screening task introduced in (Carvajal-Patiño et al., 2025). Their model is trained to approximate normalized molecular docking scores. Trained models can then be used to rank compounds by their binding likelihood to target sites, hence achieving virtual screening.
(Processing): The dataset is reproduced from the RNAmigos2 (Carvajal-Patiño et al., 2025) paper. The authors curated a list of binding sites in a similar fashion to the one described in the RNA-Site task and clustered the sites using RMScore (Zheng et al., 2019). All binders found for each cluster are retained as positive examples, and a set of drug-like chemical decoys are added as negative partners. Molecular docking scores are computed with rDock (Ruiz-Carmona et al., 2014) on all binding site-small molecule pairs. The metric used is the AuROC.
5 Implementation
Next, we briefly showcase the use of our framework for a simple programmatic access to proposed tasks. In Figure 1, we show how practitioners can access our datasets from Python code, automatically downloading them from Zenodo, choosing a representation (in this example a Pytorch Geometric graph) and directly accessing the data in a simple and reproducible fashion. Additionally, by passing a single flag to our loaders, users can choose to execute all processing logic described here to build to build each task from scratch (end-to-end reproducibility).
1from rnaglib.tasks import get_task
2from rnaglib.representations import GraphRepresentation
3
4task = get_task(task_id=’rna_site’,root=’example’)
5task.add_representation(GraphRepresentation(framework="pyg"))
6task.get_split_loaders()
7
8for batch in task.train_dataloader:
9 rna_graph = batch["graph"]
10 target = rna_graph.y
Figure 1: Obtaining a machine learning-ready split dataset only requires a few lines of code.
Due to the rapid advances in the field, we can expect that additional interesting challenges will arise in the near future, complementary to the seven tasks introduced here. Thanks to the modularity of our tool, additional tasks on RNA can be easily integrated in our framework for future releases. This is illustrated in the documentation at rnaglib.org.
6 Experiments
We hereby explicitly state that we limit the scope of this work to the creation of a useful deep learning library for RNA 3D structure based modeling. We refrain from suggesting novel model architectures; models discussed in this work uniquely serve the goal of illustrating the practical applicability of our library. Within these constraints, we highlight the utility of our benchmark by training a simple relational graph convolutional networks (RGCN) on our proposed tasks. The performance of this baseline is reported in Table 3. Additional details on the experimental setup and results across further metrics can be found in supplementary Section B. These results provide an initial baseline for future model developments on the introduced tasks.
Task Metric Score
RNA_Ligand AuROC 0.6751
RNA_CM Balanced Accuracy 0.6615
RNA_Site Balanced Accuracy 0.6309
RNA_Prot Balanced Accuracy 0.6254
RNA_IF Sequence Recovery 0.3523
RNA_VS AuROC 0.8550
RNA_GO Jaccard 0.3167
Table 3: Selected representative performance metric for each task.
Furthermore, we compare our RGCN approach to existing methods on the TR60/TE18 dataset split (Su et al., 2021) for the RNA-Site task and the dataset and splits from Joshi et al. (2024) for the RNA-VS task. Performance metrics are provided in Supplementary Tables A.2 and A.3. Despite using a relatively simple model, our approach achieves competitive results, aligning with recent methods while falling short of state-of-the-art performance.
7 Discussion
We have introduced a versatile and modular library designed to advance the application of deep learning to RNA structural analysis. By providing a suite of diverse and well-defined datasets and splits, our library enables robust benchmarking and facilitates the development of innovative computational models. It also ensures reproducibility and promotes standardized evaluation, fostering confidence in computational findings.
In the future, the development of additional tasks, such as the assessment of structural models (Townshend et al., 2021b) and a greater focus on RNA embeddings, presents exciting opportunities.
As the understanding of the dynamic nature of biological macromolecules evolves, so may the preferred representation for RNA. Our tool’s built-in 2.5D-graph representations offer extensive adaptability and have been used to develop recent competitive RNA structure-based models (Wang et al., 2024). Nevertheless, the library can easily be extended to incorporate new or improved representations as they emerge.
In this work, we unify various biological challenges where RNA structure is a key player, and provide direct access to high quality data and benchmarking functionality for deep learning model development and computational biology communities alike. Furthermore, the choice of integrating the benchmark in the rnaglib toolkit opens access to non-specialist researchers, and provides robust guarantees for quality and reproducibility. We hope that this benchmark will facilitate the development of new deep learning tools in the highly promising area of RNA structural biology.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
Appendix A Details about datasets
In this appendix, we first detail more extensively the datasets and data preprocessing applied (Section A). Then, we provide additional results of the models tested on our tasks together with details about training and hyperparameters used to perform the experiments (Section B). Eventually, we provide the code defining RNA-CM task, which provides a model to introduce new tasks (Section LABEL:sup:sec:code).
A.1 Datasets and general preprocessing
In this section, we provide additional insights about the data and the way they are being preprocessed.
As mentioned in section 3, RNA molecules present in the PDB files exhibit a bimodal distribution over the number of residues. Figure 1(a) displays the distribution of the RNA sizes (defined as the number of residues) of the RNAs from the PDB.
(a) Whole RNAs
(b) RNA chunks
(c) RNA chunks after size filtering
Figure A.1: Distribution of RNAs or RNA chunks sizes
In order to work at a more biologically relevant scale, we partition the raw RNAs from PDB files into connected components in order to use these connected components as samples in our different machine learning tasks (except RNA-Ligand in which the data are partitioned in a different way, into binding pockets rather than into connected components). Figure 1(c) represents the distribution of the number of residues by connected component chunks.
We remove the RNA components of insufficient size for structure-based machine learning, as well as very large components which negatively impact the computational performance of the data loading as well as model training. To this end, we filter RNA chunks and only keep the ones having between 15 and 300 nucleotides. Figure 1(c) displays the distribution of RNA chunks sizes after filtering.
The redundancy removal process we apply is the following. First, we perform sequence-based clustering based on a similarity threshold above which RNA fragments are clustered together (relying on sequence similarity metric CD-HIT (Fu et al., 2012)). Then, within each cluster, we select the sample with the highest resolution. Then we perform structure-based clustering and structure-based redundancy removal (relying on structural similarity metric US-align (Zhang et al., 2022a)) following the same procedure. Afterwards, when instantiating the splitters, we perform structure-based clustering with a different threshold to define the clusters which will be required to be grouped either in train, val or test set (which we name ”splitting clustering”). Since we only select one representative sample per cluster, the number of clusters prefigures the number of samples of the final dataset. Therefore, a tradeoff is to be considered between having a large amount of data and discarding redundant RNAs.
For this reason, we study the impact of both the redundancy removal and the splitting clustering threshold on the number of clusters generated in the case of RNA-CM task. Results are reported in figure A.2. Figure 2(a) displays, for 4 different sequence-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold. Here, the redundancy removal thresholds 0.90, 0.80 and 0.70 give the same plot since CD-HIT similarity values are strongly concentrated around 0.5 and 1. We finally choose 0.90. Figure 2(b) displays, for 4 different structure-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold. In all cases, the structure-based redundancy removal is performed after a sequence-based redundancy removal using a 0.90 threshold. In all the experiments, the splitting clustering is performed based on US-align similarity (structural similarity).
(a) Using sequence-based redundancy removal
(b) Using structure-based redundancy removal
Figure A.2: Number of clusters after similarity clustering based on the splitting clustering threshold
A.2 Task-specific data preprocessing
In this section, we provide additional details about the tasks involving a specific data preparation process.
A.2.1 RNA-Ligand
In this task, we selected the three most frequent ligands which were neither modified RNA or DNA residues nor modified protein residues, following the ligand definition proposed by Hariboss (Panei et al., 2022). We only retained the binding pockets binding to one of these three ligands in order to ensure the amount of binding pockets binding to each ligand would be significant enough to enable learning in a multi-class classification framework. The three ligands retained are paromomycin (called PAR in the PDB nomenclature), gentamycin C1A (LLL) and aminoglycoside TC007 (8UZ). Their structures are displayed in Figure A.3.
(a) Paromycin (PAR)
(b) Gentamycin C1A (LLL)
(c) Aminoglycoside TC007 (8UZ)
Figure A.3: Structures of the ligands selected for RNA-Ligand task
A.2.2 RNA-GO
When building RNA-GO, we explore all the GO-terms of RNAs from the PDB, remove the GO-terms which occur more than 1000 times (these are in fact the GO-terms corresponding to ”structural constituent of ribosome”, ”ribosome” and ”tRNA”). We also remove the GO-terms which are underrepresented (those which occur less than 50 times in our experiments). We then remove the GO-terms which are very correlated to other ones by performing GO-terms clustering based on correlation matrix (with correlation threshold 0.9) and keeping only one representative GO-term per cluster. After this preprocessing, 5 GO-terms are remaining: 0000353 (formation of quadruple SL/U4/U5/U6 snRNP), 0005682 (U5 snRNP), 0005686 (U2 snRNP), 0005688 (U6 snRNP) and 0010468 (regulation of gene expression).
Appendix B Additional results
Table A.1: Test performance metrics for various RNA-related tasks.
Task Test F1-Score Test AUC Test Global Balanced Accuracy Test MCC Test Jaccard
RNA_Ligand 0.2771 0.6751 0.4678
RNA_CM 0.1957 0.7393 0.6615 0.1695
RNA_Site 0.3346 0.5929 0.6309 0.3098
RNA_Prot 0.4545 0.6654 0.6254 0.2469
RNA_IF 0.3326 0.6201 0.3523 0.1319
RNA_VS 0.855
RNA_GO 0.4074 0.8406 0.7067 0.3167
Hyperparameters used:
RNA_Ligand: n_layers=4, hidden_dim=128, lr=0.00001, dropout=0.5
RNA_CM: n_layers=3, hidden_dim=128, lr=0.001, dropout=0.5
RNA_Site: n_layers=4, hidden_dim=256, lr=0.001, dropout=0.5
RNA_Prot: n_layers=4, hidden_dim=64, lr=0.01, dropout=0.2
RNA_IF: n_layers=3, hidden_dim=128, lr=0.0001, dropout=0.5
RNA_VS: n_layers=3, hidden_dim=64/32, lr=0.001, dropout=0.2
RNA_GO: n_layers=3, hidden_dim=64, lr=0.001, dropout=0.5
For RNA_IF, the reported value in “Global Balanced Accuracy” is sequence recovery.
Table A.2: We compare a standard RGCN using the rnaglib’s task module with various published results using the TR60/TE18 split. Note: Binding site definitions may vary slightly between models.
Methods MCC AUC
Rsite2 (Zeng & Cui, 2016) 0.010 0.474
Rsite (Zeng et al., 2015) 0.055 0.496
RBind (Wang et al., 2018) 0.141 0.540
RNAsite_seq (Su et al., 2021) 0.160 0.641
RNAsite_str (Su et al., 2021) 0.185 0.695
RNAsite (Su et al., 2021) 0.186 0.703
rnaglib RNA-Site 0.113 0.606
Table A.3: Sequence recovery scores for RNA inverse folding models. We use a standard two layer RGCN part of rnaglib’s task module on the dataset and split published by Joshi et al. (2024)
. Method Sequence Recovery gRNAde (Joshi et al., 2024) 0.568 Rosetta (Leman et al., 2020b) 0.450 RDesign (Tan et al., 2023) 0.430 FARNA (Alam et al., 2017) 0.321 ViennaRNA (Lorenz et al., 2011) 0.269 rnaglib RNA-IF 0.410

### v2

@@
 Abstract
-The RNA structure-function relationship has recently garnered significant attention within the deep learning community, promising to grow in importance as nucleic acid structure models advance.
-However, the absence of standardized and accessible benchmarks for deep learning on RNA 3D structures has impeded the development of models for RNA functional characteristics.
-In this work, we introduce a set of seven benchmarking datasets for RNA structure-function prediction, designed to address this gap.
-Our library builds on the established Python library rnaglib, and offers easy data distribution and encoding, splitters and evaluation methods, providing a convenient all-in-one framework for comparing models.
-Datasets are implemented in a fully modular and reproducible manner, facilitating for community contributions and customization.
-Finally, we provide initial baseline results for all tasks using a graph neural network.
-RNA Structure-Function Relationship, Graph Learning, Benchmarks, RNA 3D Structure, Computational Biology, RNA Structural Modeling
-Source code: github.com/cgoliver/rnaglib
-Documentation: rnaglib.org
+The relationship between RNA structure and function has recently attracted interest within the deep learning community, a trend expected to intensify as nucleic acid structure models advance.
+Despite this momentum, a lack of standardized, accessible benchmarks for applying deep learning to RNA 3D structures hinders progress.
+To this end, we introduce a collection of seven benchmarking datasets specifically designed to support RNA structure–function prediction.
+Built on top of the established Python package rnaglib, our library streamlines data distribution and encoding, provides tools for dataset splitting and evaluation, and offers a comprehensive, user-friendly environment for model comparison.
+The modular and reproducible design of our datasets encourages community contributions and enables rapid customization.
+To demonstrate the utility of our benchmarks, we report baseline results for all tasks using a relational graph neural network.
+Data and code are available at github.com/cgoliver/rnaglib.
 1 Introduction
-Recent years have witnessed the advent of deep learning methods for structural biology culminating in the award of the Nobel Prize in Chemistry.
+Recent years have witnessed the advent of deep learning methods for structural biology culminating in the award of the Nobel Prize in Chemistry in 2024.
 AlphaFold (Jumper et al., 2021) revolutionized protein structure prediction, equipping the field with millions of new structures.
 Breakthroughs go beyond structure prediction, notably in protein design (Watson et al., 2023; Dauparas et al., 2022), drug discovery (Schneuing et al., 2024; Corso et al., 2022) or fundamental biology (van Kempen et al., 2022).
-While it is tempting to attribute the success of these methods to the increase in available structural data caused by AlphaFold, most of the methods are actually not reliant on them.
-Instead, it seems that these breakthroughs result from progress in training neural encoders that directly model protein structures (Jing et al., 2020; Zhang et al., 2022b; Gainza et al., 2020; Wang et al., 2022).
-This progress is in turn rooted in solid competitions (CASP, CAPRI), and benchmarks (Townshend et al., 2021a; Kucera et al., 2023; Zhu et al., 2022; Jamasb et al., 2024; Notin et al., 2023).
-By setting clear goals, such benchmarks are the foundation for the development of structure encoders.
-Yet to date, structure-function benchmarks have focused on proteins.
-Ribonucleic acids (RNAs) are a large family of molecules which support biological functions along every branch of the tree of life.
-Besides messenger RNAs, non-coding RNAs carry out biological functions by adopting complex 3D folds (Cech & Steitz, 2014) like proteins do and take up diverse roles in cellular functions, including gene regulation, RNA processing, and protein synthesis (Statello et al., 2021).
+A key factor in these breakthroughs was the development of neural encoders that directly model protein structure (Jing et al., 2020; Zhang et al., 2022b; Gainza et al., 2020; Wang et al., 2022).
+These advances in turn are rooted in solid competitions (CASP, CAPRI), and benchmarks (Townshend et al., 2021a; Kucera et al., 2023; Zhu et al., 2022; Jamasb et al., 2024; Notin et al., 2023).
+By setting clear goals, such benchmarks are the foundation for the development of structure encoders and are required to leverage ever increasing structural data availability.
+Yet, to date, structure-function benchmarks have focused on proteins, hampering model development for other biological macromolecules.
+Ribonucleic acids (RNAs) are a large family among these macromolecules, supporting biological functions along every branch in the tree of life.
+Besides messenger RNAs which code for proteins, non-coding RNAs are a vast class of molecules which can adopt complex 3D folds (Cech and Steitz, 2014) and take up diverse roles in cellular functions, including gene regulation, RNA processing, and protein synthesis (Statello et al., 2021).
 However, our understanding of non-coding RNAs and their functions remains limited.
@@
 Another predominant challenge to a functional understanding of RNA 3D structure lies in the lack of infrastructure for the development and evaluation of function prediction models.
-In this work, we propose a benchmarking suite to act as this facilitating framework.
+In this work, we propose a benchmarking suite to act as a facilitating library for the development of structure encoders on RNA.
 Our key contributions include:
@@
 Each task consists of a dataset, a splitting strategy, and an evaluation method, laying the ground for comparable, reproducible model development.
-  • End-to-end reproducible and modular access to task data.
-Modular annotators, filters and splitting strategies, both novel and from existing literature, facilitate the addition of new tasks by other researchers across fields.
+  • Modular annotators, filters and splitting strategies, both novel and from existing literature, that facilitate the construction of new and custom tasks by other researchers, reducing the time spent on data preparation while ensuring reproducibility of achieved results.
+  • We apply the benchmark to simple architectures to initialize a leaderboard and investigate the effect of splitting and structural context on performance.
 2 Related Work
 2.1 Protein Focused Benchmarking
-Classic tasks on proteins appeared independently in unrelated papers, such as GO term and EC number prediction (Gligorijević et al., 2021), fold classification (Hou et al., 2018), binding site detection and classification (Gainza et al., 2020) or binding affinity regression (Wang et al., 2005a).
+Classic tasks on proteins appeared independently in unrelated papers, such as GO term and EC number prediction (Gligorijević et al., 2021), fold classification (Hou et al., 2018), binding site detection and classification (Gainza et al., 2020) or binding affinity regression (Wang et al., 2005).
 To our knowledge, ATOM3D (Townshend et al., 2021a) was the first systematic benchmark for molecular systems, albeit heavily focused on proteins.
@@
 2.2 RNA Structural Datasets and Benchmarking
-In the realm of RNA 3D structure based modeling infrastructure, three papers propose cleaned datasets with the objective of facilitating machine learning.
-RNANet (Becquey et al., 2021) proposes a dataset joining RNA structures with their corresponding sequence alignment.
-RNAsolo (Adamczyk et al., 2022) provides access to cleaned RNA files in various formats through a web interface.
-Finally, RNA3DB (Szikszai et al., 2024) offers a curated dataset specifically designed for RNA structure prediction models.
-None of these methods propose benchmark tasks to compare RNA modeling and learning methods.
-Moreover, some datasets along with splits were independently proposed in several work, pertaining to small-molecules binding sites (Wang et al., 2018), or virtual screening (Panei et al., 2022; Carvajal-Patiño et al., 2025).
-The inverse folding field has seen a particular breath of advances in (Joshi et al., 2024; Tan et al., 2023, 2025; Huang et al., 2024; Wong et al., 2024).
-Recent benchmarking suites like BEACON (Ren et al., 2024) provide extensive tools for models on RNA, but do not rely on structural data or focus on non-coding RNAs.
+In the realm of RNA 3D structure based modeling infrastructure, three papers propose datasets with the objective of facilitating machine learning.
+They either provide cleaned files in various formats through a web interface (RNAsolo (Adamczyk et al., 2022)), join RNA structures with their corresponding sequence alignment (RNANet (Becquey et al., 2021)) or focus on structure prediction (RNA3DB (Szikszai et al., 2024)).
+Recently, multiple benchmarking suites were introduced for models on RNA sequences (BEACON (Ren et al., 2024)), predicted structures of short RNAs (Xu et al., 2024), or for fitness prediction (Arora et al., 2025).
+None of these methods propose benchmark tasks to compare RNA modeling and learning methods on experimental structures.
 2.3 Structure-Based RNA Models
@@
 RNAmigos and RNAmigos2 (Oliver et al., 2020; Carvajal-Patiño et al., 2025) proposed incorporating non-canonical interactions in the graph, in conjunction with metric learning pretraining.
-In a follow-up work (Wang et al., 2023), a similar method for binding site prediction was used.
-Finally, gRNAde (Joshi et al., 2024) adapted the popular GVP (Jing et al., 2021) protein encoder, which uses the atomic position in the message passing algorithm, to RNA.
-Despite manifold efforts at deep learning-based modeling RNA 3D structure, we lack the means to systematically compare models and onboard new practitioners.
+ARES (Townshend et al., 2021b) applied a tensor field neural network (Thomas et al., 2018) to RNA structure, modeling it as an atomic graph.
+Geometric Vector Perceptron (Jing et al., 2021), an equivariant message passing protein encoder, was adapted to RNA in (Joshi et al., 2024; Tan et al., 2023, 2025; Huang et al., 2024; Wong et al., 2024).
+EquiRNA (Li et al., 2025) proposed a hierarchical architecture.
+Lately, HARMONY (Xu et al., 2025) introduces a model using multiple RNA modalities.
+Despite booming interest in deep learning-based modeling RNA 3D structure, we lack the means to systematically compare models and onboard new practitioners.
 3 Tools to assemble tasks
-This section describes the full process of building a task: data collection, quality filtering, cleanup, and splitting.
-Table 1: Description of the proposed datasets and biologically-relevant tasks.
-Task Name Short description Novel Expanded
-1. RNA-GO Classify RNAs into one or more frequent function. $\surd$
-2. RNA-IF Find a sequence that folds into this structure. $\surd$
-3. RNA-CM Predict which RNA residues are chemically modified from its geometry. $\surd$
-4. RNA-Prot Predict which RNA residues are interacting with a protein. $\surd$
-5. RNA-Site Predict which RNA residues are interacting with a small-molecule protein. $\surd$
-6. RNA-Ligand Classify an RNA binding site, based on the ligand it binds to. $\surd$
-7. RNA-VS Predict the affinity of a small molecule for an RNA binding site, and use it for Virtual Screening (VS). $\surd$
+We describe the full process of building a task: data collection, quality filtering, cleanup, and splitting.
+We developed these tools to build our proposed tasks.
+They are ready to use and documented for any practitioner wishing to build his own reproducible task with minimal effort.
 Data collection and annotation
 Our data originates from the RCSB-PDB DataBank (Kouranov et al., 2006), where we fetch all 3D structures containing RNA.
-In some tasks, filter down the set of RNA-containing PDB structures with the subset proposed in (Leontis & Zirbel, 2012) and later referred to as bgsu.
+In some tasks, filter down the set of RNA-containing PDB structures with the subset proposed in (Leontis and Zirbel, 2012) and later referred to as bgsu.
 We annotate each system with RNA-level features such as its resolution, and residue-level features such as their coordinates, the presence of interacting compounds, and the amount of protein atoms in the residues’ vicinities.
 Structure partitioning and quality filters
-RNA molecules present in PDB files exhibit a bimodal distribution over the number of residues (see figures 1(a) and 1(b)).
-Many systems have less than 300 residues, while a few others (mostly ribosomal structures) have several thousands.
-To avoid discarding entire systems, we split whole systems into sub-structures by sets of connected components in the graph formed by all base pairings and backbone connections.
-Concretely, this allows independent RNA fragments appearing in the same PDB file to be treated as multiple systems; the resulting decrease in system size leads to a reduction in required computing power for downstream steps, such as similarity-based splitting.
+RNA molecules present in PDB files exhibit a bimodal distribution over the number of residues (see figures 5(a) and 5(b)).
+Many systems have less than 300 residues, while some (mostly rRNAs) have several thousands.
+We consider independent RNA fragments from the same PDB file should be treated as multiple systems ; we split each of our systems into connected components in the graph formed by RNA residues and their interactions.
+Concretely, the resulting decrease in system size leads to a reduction in required computing power for downstream steps.
 We implement resolution and size filters, with default cutoff values including systems below 4Å resolution and with a length of 15 to 500 residues.
-In tasks, the upper size cutoff is used to limit computational expenses.
+In tasks, the upper size cutoff is used to limit computational expense.
 We also provide a protein content filter which removes RNA structures that are heavily structured through an interaction with a protein.
-This crucial filter has been overlooked in most of existing RNA structural datasets.
+This crucial filter has been overlooked in most existing RNA structural datasets.
 In addition, we implement the drug-like filter introduced in Hariboss (Panei et al., 2022) for small molecules binding to RNA.
-Minimizing data leaks
-Because biological data points are often related in terms of evolutionary history and topological characteristics, insufficient rigor in splitting can lead to severe data leaks.
-Unsurprisingly, numerous examples of data leakage have pestered structural biology learning methods, one famous example being the first PDBbind dataset that was shown to incorporate several severe biases (Wang et al., 2005b; Volkov et al., 2022).
-Efforts like the PINDER database (Kovtun et al., 2024) have set a standard in the field for model generalisation assessment.
-In our work, we implement the computation of sequence-based or structure-based similarity matrices, with CD-HIT (Fu et al., 2012) and US-align (Zhang et al., 2022a) respectively.
-Given a similarity matrix $S$, computed over a set of RNA molecules $\mathcal{R}$, and a threshold $\theta$, we introduce the matrix $S^{\theta}$ defined as $S^{\theta}_{i,j}=\delta\{S_{i,j}\geq\theta\}$, where $\delta$ is the indicator function.
-We propose a clustering algorithm that considers $(\mathcal{R},S^{\theta})$ as a graph, and returns connected components as clusters.
+Redundancy filtering and dataset splitting
+Biological data points are often related in terms of evolutionary history and topological characteristics.
+Moreover, as structures of a given system are often observed in slightly different conditions, it is common to obtain copies with minimal variations, introducing redundancy in the data.
+Insufficient rigor in splitting can thus lead to severe data leaks (Bernett et al., 2024; Volkov et al., 2022).
+Efforts like the PINDER database (Kovtun et al., 2024) have set a standard in the field for model generalization assessment.
+Given a set of RNA molecules $\mathcal{R}$, a similarity metric ${s}$ and a threshold $\theta$, denote as $\mathcal{G}$ the graph $(\mathcal{R},\mathcal{E}^{\theta})$, where $\mathcal{E}^{\theta}:=\{(r_{1},r_{2})\in\mathcal{R}^{2},s(r_{1},r_{2})>\theta\}$.
+We propose a clustering algorithm that returns connected components of $\mathcal{G}$ as clusters.
 This ensures that any pair of points in different clusters has a maximal similarity of $\theta$.
-Redundancy filtering and dataset splitting
-Starting with clusters, we propose a redundancy removal algorithm that selects the element with the highest resolution for each cluster.
-In most tasks, we first apply this algorithm at a sequence similarity threshold of 0.9.
-This stringent threshold discards copies of systems with minimal variations that are common when getting the structure of a given system in slightly different conditions.
-Then, we apply another structure-based redundancy threshold of 0.8.
-In addition, we propose a novel splitting algorithm that starts from clustered data, but uses a less conservative structural similarity threshold (0.5 unless mentioned otherwise).
-Then, we aggregate those clusters together to form splits of a certain size, optionally following a label balancing secondary objective, using the linear programming PuLP (Mitchell et al., 2011).
-This prevents data leakage between splits, while allowing split stratification based on labels.
+In our work, we compute sequence-based with CD-HIT (Fu et al., 2012) or structure-based similarity with US-align (Zhang et al., 2022a).
+We propose a redundancy removal algorithm that selects the highest resolution system for each cluster.
+For most tasks, we first apply this algorithm at a sequence similarity threshold of 0.9, and then at a structure similarity threshold of 0.8.
+rnaglib includes a data splitting algorithm that aggregates clusters together.
+To prevent data leakage, we build clusters with a more conservative structural similarity threshold (0.5 unless mentioned otherwise).
+Then, we cast the aggregation of clusters as an integer programming problem, solved using the linear programming tool PuLP (Mitchell et al., 2011).
+This enforces structural dissimilarity between splits for predetermined split size ratios while achieving optimal label stratification.
 Metrics computation
@@
 binary classification, multi-label, etc.), each task implements a standardized way to compute evaluation metrics.
-These are specified for each task in Section 4.
+Each task is evaluated by a recommended metric, specified in Table 2, and a range of additional standard metrics.
 Anatomy of a task
-A “task” is now simply the collection of the aformentioned components with the choices relevant to a particular biological problem.
-That is, each task consists of (i) a dataset drawn from our 3D structure database and processed using our annotations, partitions and filters, with redundancy removing (ii) fixed splits based on sequence, structure or existing literature, and (iii) a well-defined evaluation protocol.
+A “task” is now simply the collection of the aforementioned components with the choices relevant to a particular biological problem.
+That is, each task consists of (i) a dataset drawn from our 3D structure database and processed using our annotations, partitions and filters, (ii) redundancy removing and fixed splits based on sequence, structure or existing literature, and (iii) a well-defined evaluation protocol.
 This modular design will help other practitioners propose additional benchmark tasks, and lower the barrier to entry for training models on RNA structure.
 4 Tasks on RNA 3D Structure
-Our tasks suite explores various dimensions of RNA structural research briefly summarized in Table 1.
-We offer seven tasks among which three are based on previous research and offer publicly available split datasets (RNA-IF (Joshi et al., 2024), RNA-Site (Su et al., 2021) and RNA-VS (Carvajal-Patiño et al., 2025)).
-Moreover, for two of these three tasks, we propose an enhanced, expanded version, resulting in a total of nine separate datasets.
+Our tasks suite explores various dimensions of RNA structural research briefly summarized and illustrated in Figure 1.
+We offer seven tasks among which three are based on previous research with published datasets (RNA-IF (Joshi et al., 2024), RNA-Site (Su et al., 2021) and RNA-VS (Carvajal-Patiño et al., 2025)).
+For RNA-IF and RNA-Site, we propose an enhanced, expanded version, resulting in a total of nine separate datasets.
 Armed with an RNA structure, the main focus of each of the subsequent tasks is to either predict aspects of the RNA’s biological function, or to reverse the direction and predict the sequence from which a given structure arises (RNA design).
 We provide an overview of these tasks from a machine learning point of view relating to the type of task and dataset size in Table 1.
-Table 2: Description of the ML features of each of our tasks.
-The number in parenthesis refers to the number of class, in the multi-class setting.
-We include a comparison of the datasets we introduce with existing ones.
+Task Name Short description
+1. RNA-GO Classify RNAs into one or more frequent function.
+2. RNA-IF Find a sequence that folds into this structure.
+3. RNA-CM Predict which RNA residues are chemically modified from its geometry.
+4. RNA-Prot Predict which RNA residues are interacting with a protein.
+5. RNA-Site Predict which RNA residues are interacting with a small-molecule protein.
+6. RNA-Ligand Classify an RNA binding site, based on the ligand it binds to.
+7. RNA-VS Predict the affinity of a small molecule for an RNA binding site, and use it for Virtual Screening (VS).
+Figure 1: A graphical overview of the seven tasks included in rnaglib illustrated through a precursor tRNA (4GCW) and a SAM-I riboswitch (5FJC).
+The integrated table contains a short summary of each task.
+4.1 RNA-GO: Function Tagging
+(Definition): This is an RNA-level, multi-label classification task where an RNA is mapped to five possible labels representing molecular functions.
+(Context): Effective function prediction models have the capacity to uncover new structure-function connections.
+The Gene Ontology (GO) (Ashburner et al., 2000) was developed to associate a function to a gene, and thus, indirectly to the RNA or protein it encodes.
+It resulted in discrete functional categories called GO terms.
+Predicting the GO term from a protein structure was proposed as a task in (Gligorijević et al., 2021) and has since been regularly used as a benchmark task.
+A manual annotation of RNA sequences with GO terms is available in Rfam (Griffiths-Jones et al., 2003; Ontiveros-Palacios et al., 2025), a database of non-coding RNA families.
+(Processing): We extracted all GO terms corresponding to an RNA structure.
+Then, we filtered these terms based on their frequency, removing the two most frequent labels (ribosomal systems and tRNAs), as well as infrequent ones (less than 50 occurrences).
+We then grouped together GO-terms that were fully correlated, which resulted in five classes.
+Finally, we discarded RNA fragments with less than 15 residues, resulting in 501 systems.
+We split these systems based on a sequence similarity cutoff, following Gligorijević et al.
+(2021).
+4.2 RNA-IF: Molecular Design
+(Definition): This is a residue-level, classification task. Given an RNA structure with masked sequence information (keeping only the coordinates of the backbone), we aim to recover the sequence.
+(Context): One avenue for designing new macromolecules is to design a coarse grained tertiary structure first, for instance scaffolding a structural motif, and then design a sequence that will fold into this given structure.
+This second step is denoted as Inverse Folding (IF), since it maps a known structure to an unknown sequence.
+Inverse folding is a well-established task in protein research with many recent breakthroughs (Watson et al., 2023), with adaptations for RNA with classical models (Leman et al., 2020a).
+Learning based approaches were pioneered by RDesign (Tan et al., 2023), gRNAde (Joshi et al., 2024) and other papers specifically address backbone generation (Anand et al., 2024) and protein-binding RNA design (Nori and Jin, 2024).
+(Processing): We gather all connected components in our 3D structure database having sizes between 15 and 300.
+Then, we remove identical sequences using a redundancy removal step with CD-HIT with a cutoff of 0.99.
+We cluster the data on structural similarity with a threshold of 0.9, and split these while ensuring a maximal similarity of 0.5.
+We also provide datasets and splits from gRNAde (Joshi et al., 2024).
+Our dataset differs by allowing multi-chain systems, a stricter size cutoff, and a duplicates filter, leading to a reduced dataset size.
+4.3 RNA-CM: Dynamic Function Modulation
+(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict which, if any, of its residues are chemically modified.
+(Context): In addition to the four canonical nucleotides, more than 170 modified nucleotides were found to be integrated in RNA polymers (Boccaletto et al., 2022), affecting the functions of a diverse range of ncRNA (Roundtree et al., 2017).
+We propose to detect such chemical modifications from subtle perturbations of the canonical geometry of the RNA backbones.
+(Processing): To build this task, we start from the entire dataset, partition it into connected components and apply our size filter.
+Then, we filter for systems that include modified residues, relying on PDB annotations.
+We apply our default redundancy removal and splitting strategies, which results in 185 data points.
+Table 1: Description of the ML characteristics of each of our tasks.
+The number in parenthesis refers to the number of classes.
+We include published reference datasets available through rnaglib below their corresponding tasks.
 Task name Level Type Dataset split sizes
@@
 7. RNA-VS (Carvajal-Patiño et al., 2025) Pocket Regression 304-34-65
-4.1 RNA-GO: Function Tagging
-(Definition): This is a multi-label classification task where input RNA is mapped to a sequence of 5 possible labels representing molecular functions.
-(Context): Effective function prediction models have the capacity to uncover new structure-function connections.
-The Gene Ontology (GO) (Ashburner et al., 2000) was developed to associate a function to a gene, and thus, indirectly to the RNA or protein it encodes.
-It resulted in discrete functional categories called GO terms.
-Predicting the GO term from a protein structure was proposed as a task in (Gligorijević et al., 2021) and has since been regularly used as a benchmark task.
-A mapping between RNA sequences and GO terms is available in Rfam (Griffiths-Jones et al., 2003; Ontiveros-Palacios et al., 2025), a database of non-coding RNA families which are manually annotated with GO terms.
-(Processing): We extracted all GO terms corresponding to an RNA structure.
-Then, we filtered these terms based on their frequency, removing the two most frequent labels (ribosomal systems and tRNAs), as well as infrequent ones (less than 50 occurrences).
-We then grouped together GO-terms that were fully correlated, which resulted in five classes.
-Finally, we discarded RNA fragments with less than 15 residues, resulting in 501 systems.
-We split these systems based on a sequence similarity cutoff, following Gligorijević et al.
-(2021).
-Accuracy, AuROC and the Jaccard index are the metrics computed for evaluation.
-4.2 RNA-IF: Molecular Design
-(Definition): This is a residue-level, classification task where given an RNA structure, we mask all residue identities (by keeping only the coordinates of the backbone) and aim to learn the mapping from structure to sequence.
-(Context): One avenue for designing new macromolecules is to design a coarse grained tertiary structure first, for instance scaffolding a structural motif, and then design a sequence that will fold into this given structure.
-This second step is denoted as Inverse Folding (IF), since it maps a known structure to an unknown sequence.
-Inverse folding is a well-established task in protein research with many recent breakthroughs (Watson et al., 2023).
-It has recently been adapted for RNA with classical models such as Rosetta (Leman et al., 2020a).
-Learning based approaches were pioneered by gRNAde (Leman et al., 2020a; Joshi et al., 2024) and other papers specifically address backbone generation (Anand et al., 2024) and protein-binding RNA design (Nori & Jin, 2024).
-(Processing): We gather all connected components in our 3D structure database with a size between 15 and 300.
-Then, we removed identical sequences using a redundancy removal step with CD-HIT with a cutoff of 0.99.
-We cluster the data on structural similarity with a threshold of 0.9, and split these while ensuring a maximal similarity of 0.5.
-We also provide datasets and splits from gRNAde (Joshi et al., 2024), that were obtained in a similar manner.
-Our dataset differs by allowing multi-chain systems, a stricter size cutoff, and a duplicates filter, leading to a reduced dataset size.
-This task is evaluated with sequence recovery.
-4.3 RNA-CM: Dynamic function modulation.
-(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict which, if any, of its residues are chemically modified.
-(Context): In addition to the four canonical nucleotides, more than 170 modified nucleotides were found to be integrated in RNA polymers (Boccaletto et al., 2022).
-Multiple functions of a diverse range of ncRNA have been shown to be directly dependent on such chemical modifications (Roundtree et al., 2017).
-We propose to detect such chemical modifications from subtle perturbations of the canonical geometry of the RNA backbones.
-Training a predictor on RNA structures is not the most straightforward way of detecting chemical modifications, but may uncover function-modulating modifications that hitherto went unnoticed.
-(Processing): To build this task, we start from the whole dataset, partition it into connected components and apply our size filter.
-Then, we filter for systems that include modified residues, relying on PDB annotations that flag such modified residues.
-We apply our default redundancy removal and splitting strategies, which results in 185 data points.
-Performance is evaluated with accuracy and AuROC.
-4.4 RNA-Prot: Biological complex modeling.
+4.4 RNA-Prot: Biological Complex Modeling
 (Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether a protein residue is closer than 8Å to any of its residues.
@@
 RNA structure is often, but not always, heavily disrupted upon interaction with a protein.
-We expect this task to be an easier version of the RNA-CM task.
 (Processing): To build this task, we start from the bgsu non-redundant dataset, partition it into connected components and apply our size and resolution filters.
 We remove systems originating from ribosomes that are fully encapsulated in a complex.
-Then, we retain only RNA interacting with a protein and apply our default redundancy removal and splitting strategies, resulting in 1251 data points.
-Performance is evaluated with accuracy and AuROC.
+Then, we retain only RNA interacting with a protein and apply our default redundancy removal and splitting strategies, yielding 1251 data points.
 Tasks for RNA Drug Design
-Further tasks aim to address RNA small-molecule drug-design.
-RNA is increasingly recognized as a promising family of targets for the development of novel small molecule therapeutics  (Falese et al., 2021; Haga & Phinney, 2023; Disney, 2019; Abulwerdi et al., 2019).
+RNA is increasingly recognized as a promising family of targets for the development of novel small molecule therapeutics  (Falese et al., 2021; Haga and Phinney, 2023; Disney, 2019; Abulwerdi et al., 2019).
 Being able to target RNA would drastically increase the size of the druggable space, and propose an alternative for overused protein targets in scenarios where they are insufficient.
 In addition, they represent a therapeutic avenue in pathologies where protein targets are absent, such as in triple-negative breast cancer (Xu et al., 2020).
-4.5 RNA-Site: Drug target detection.
+The following tasks aim to address questions in RNA small-molecule drug-design.
+4.5 RNA-Site: Drug Target Detection
 (Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether a ligand is closer than 6Å to any of its residues.
@@
 We start by following similar steps as for RNA-Prot (without removal of ribosomal systems).
-We then include a drug-like filter on the small-molecule side.
+We then include a drug-like filter on the small-molecules.
 Finally, we include a protein-content filter, removing systems with more than 10 protein atom neighbors, ensuring that the binding is modulated by RNA only.
 We use the default redundancy removal and splitting, resulting in 458 systems.
-The predictions are evaluated using accuracy and AuROC.
 4.6 RNA-Ligand: Pocket Categorization
@@
 (Processing): Starting with similar steps as the RNA-Site, we obtain a set of structures that display RNAs in interaction with one or more drug-like small molecules.
-We then proceed to extracting the context of the binding pocket by seeding two rounds of breadth-first search with all residues closer than 8Å to an atom of the binder.
+We then proceed to extract the context of the binding pocket by seeding two rounds of breadth-first search with all residues closer than 8Å to an atom of the binder.
 This results in all binding pockets in our database and we group these pockets with a sequence clustering.
@@
 We then compute the most frequent ligands among this dataset and retain only the top three.
-Subsequently, we discard all binding events to other small molecules, and add clusters that bind only one of the top three ligands to our dataset.
-This dataset is split on structural similarity and evaluated on MCC and AuROC.
+Subsequently, we discard all binding events to other small molecules, and add clusters that bind only one of the top three ligands to our dataset and split based on structural similarity.
 4.7 RNA-VS: Drug Screening
@@
 Molecular docking scores are computed with rDock (Ruiz-Carmona et al., 2014) on all binding site-small molecule pairs.
-The metric used is the AuROC.
 5 Implementation
@@
 In Figure 2, we show how practitioners can access our datasets from Python code, automatically downloading them from Zenodo, choosing a representation (in this example a Pytorch Geometric graph) and directly accessing the data in a simple and reproducible fashion.
-Additionally, by passing a single flag to our loaders, users can choose to execute all processing logic described here to build to build each task from scratch (end-to-end reproducibility).
+Additionally, by passing a single flag (recompute=True) to the task fetcher, users can choose to execute all processing logic from scratch (end-to-end reproducibility).
 1from rnaglib.tasks import get_task
@@
 10 target = rna_graph.y
-Figure 1: Obtaining a machine learning-ready split dataset only requires a few lines of code.
+Figure 2: Obtaining a machine learning-ready split dataset requires only a few lines of code.
 Due to the rapid advances in the field, we can expect that additional interesting challenges will arise in the near future, complementary to the seven tasks introduced here.
@@
 This is illustrated in the documentation at rnaglib.org.
-6 Experiments
+6 Experiments and Results
 We hereby explicitly state that we limit the scope of this work to the creation of a useful deep learning library for RNA 3D structure based modeling.
 We refrain from suggesting novel model architectures; models discussed in this work uniquely serve the goal of illustrating the practical applicability of our library.
-Within these constraints, we highlight the utility of our benchmark by training a simple relational graph convolutional networks (RGCN) on our proposed tasks.
-The performance of this baseline is reported in Table 3.
-Additional details on the experimental setup and results across further metrics can be found in supplementary Section B.
-These results provide an initial baseline for future model developments on the introduced tasks.
-Task Metric Score
-RNA_Ligand AuROC 0.6751
-RNA_CM Balanced Accuracy 0.6615
-RNA_Site Balanced Accuracy 0.6309
-RNA_Prot Balanced Accuracy 0.6254
-RNA_IF Sequence Recovery 0.3523
-RNA_VS AuROC 0.8550
-RNA_GO Jaccard 0.3167
-Table 3: Selected representative performance metric for each task.
+However, we provide a baseline performance for each of our proposed tasks as a reference for future works and conduct some investigations highlighting rnaglib’s use in understanding and building RNA structure-function models.
+All reported results are conducted over three seeds and code is available from github.com/BorgwardtLab/rnaglib-experiments.
+6.1 Baseline Performance
+The baseline performance for each of our proposed tasks is achieved by training Relational Graph Convolutional Networks (RGCN) on graph representations of the tasks’ underlying datasets.
+We choose to use 2.5D graphs, whose nodes represent residues and edges represent both canonical and non-canonical interactions.
+This approach was previously shown to outperform a secondary structure based alternative in a drug-discovery task (Oliver et al., 2020).
+We report our results in Table 2 and additional details on the experimental setup, including hyperparameters, and further metrics can be found in the supplementary Section A.2.
+We observe that on all tasks, our machine learning models are able to capture a meaningful signal that outperforms naive approaches, always predicting the majority class, validating the relevance of our tasks for benchmarking purposes.
+Table 2: Baseline performance reported using our recommended representative metric for each task. We compare to always predicting the majority class.
+Task Metric Majority Score
+RNA-GO Jaccard 0.2 0.396 $\pm$ 0.023
+RNA-IF Sequence Recovery 0.28 0.372 $\pm$ 0.005
+RNA-CM Balanced Accuracy 0.5 0.633 $\pm$ 0.007
+RNA-Prot Balanced Accuracy 0.5 0.556 $\pm$ 0.004
+RNA-Site Balanced Accuracy 0.5 0.655 $\pm$ 0.012
+RNA-Ligand AuROC 0.33 0.650 $\pm$ 0.124
+RNA-VS AuROC 0.5 0.759 $\pm$ 0.005
 Furthermore, we compare our RGCN approach to existing methods on the TR60/TE18 dataset split (Su et al., 2021) for the RNA-Site task and the dataset and splits from Joshi et al.
 (2024) for the RNA-VS task.
-Performance metrics are provided in Supplementary Tables A.2 and A.3.
-Despite using a relatively simple model, our approach achieves competitive results, aligning with recent methods while falling short of state-of-the-art performance.
+Performance metrics are provided in Supplementary Tables 4 and 5.
+Despite using a significantly simpler model, our approach achieves competitive results, aligning with recent methods while falling short of state-of-the-art performance.
+6.2 Redundancy and Splitting
+To showcase rnaglib’s use in choosing suitable splitting approaches, we compare the three main splitting approaches it provides and present results in Figure 3(a).
+As expected, training a model on randomly split data leads to inflated results in some cases.
+However, the difference is not clear cut 3(a).
+Since our tasks’ datasets already pass a redundancy removal step, we postulate that the reduced redundancy mitigates the effect of leakage.
+Indeed, forgoing the redundancy removal for an example task (RNA-Site) markedly increases the performance difference between similarity-based and random splitting approaches.
+rnaglib permits setting different structure and sequence similarity thresholds for splitting, depending on use case and data availability.
+In Figure 3(b), we plot the performance of our baseline model on the redundant version of RNA-Site, varying the similarity threshold used.
+Our results show that more stringent splitting leads to a decreased, more realistic performance.
+The trade-off between leakage and learning signal can thereby be tweaked by practitioners.
+We point out that structure and sequence thresholds are not directly comparable, since their calculation differs.
+Setting a threshold of 1 results in isolated points for our structure similarity metric, resulting in comparable results to random splitting.
+This is in contrast with our sequence-based similarity metric, whose value is exactly one for highly similar structures.
+Using either a non-redundant dataset or rigorous splitting, alleviates the risk of data leakage, using both maximizes the reliability of the results.
+(a) Figure 3a: Redundancy removal has a dampening effect on data leakage across splitting strategies.
+(b) Figure 3b: In a redundant setting, stricter splitting thresholds reduce data leakage, shown for RNA-Site.
+6.3 Model Benchmarking
+We provide two case studies on selected tasks elucidating how rnaglib can assist in choosing modeling modalities.
+We investigate the impact of increasing the depth of our model, as well as the effect of using different representations.
+We focus on RNA-Site, RNA-CM, and RNA-Prot.
+First, we assess the impact of varying the number of message passing layers to understand the effect of varying structural context size on model performance.
+We train models of various depths (number of layers ranging from two to six) on each task of these three tasks, tuning only the learning rate which we adjust for optimal convergence, keeping the remaining hyperparameters as in Supplementary Table 3.
+Results are presented in Figure 4(a).
+While using four layers generally results in a good performance, the best value varies from one task to another.
+For example, predicting protein binding sites (RNA-Prot) might require less structural context than other tasks.
+[…]
+(a) Figure 4a: RGCN depth
+(b) Figure 4b: Structural detail in input RNAs
+Next, we assess various types of graph representations for RNA structures.
+We contrast 2.5D RNA graphs, whose edges include both canonical and non-canonical interactions, rnaglib’s default representation, with 2D RNA graphs (i.e.
+canonical interactions only).
+Within the latter, we consider a 2D setting, where all canonical interactions are encoded as one edge type, and a 2D+ setting, where canonical Watson-Crick base-pairing and backbone covalent interactions are distinguished (the backbone’s directionality results in a total of three edge types).
+We compare these three representations in Figure 4(b) and observe that including non canonical interactions results in a performance increase across all tasks.
+These results are in accordance with those presented in Oliver et al.
+(2020).
 7 Discussion
@@
 By providing a suite of diverse and well-defined datasets and splits, our library enables robust benchmarking and facilitates the development of innovative computational models.
-It also ensures reproducibility and promotes standardized evaluation, fostering confidence in computational findings.
-In the future, the development of additional tasks, such as the assessment of structural models (Townshend et al., 2021b) and a greater focus on RNA embeddings, presents exciting opportunities.
+It ensures reproducibility and promotes standardized evaluation, fostering confidence in computational findings.
+Its flexibility allows researchers to easily build additional tasks specific to their research needs in a wide range of RNA structural studies.
+In the future, the development of additional tasks, such as the assessment of structural models (Townshend et al., 2021c) and a greater focus on RNA embeddings, presents exciting opportunities.
 As the understanding of the dynamic nature of biological macromolecules evolves, so may the preferred representation for RNA.
@@
 We hope that this benchmark will facilitate the development of new deep learning tools in the highly promising area of RNA structural biology.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix
-Appendix A Details about datasets
-In this appendix, we first detail more extensively the datasets and data preprocessing applied (Section A).
-Then, we provide additional results of the models tested on our tasks together with details about training and hyperparameters used to perform the experiments (Section B).
-Eventually, we provide the code defining RNA-CM task, which provides a model to introduce new tasks (Section LABEL:sup:sec:code).
-A.1 Datasets and general preprocessing
+Appendix A Technical Appendices and Supplementary Material
+A.1 Details about datasets
+In this appendix, we first detail more extensively the datasets and data preprocessing applied (Section A.1).
+Then, we provide additional results of the models tested on our tasks together with details about training and hyperparameters used to perform the experiments (Section A.2).
+A.1.1 Datasets and general preprocessing
 In this section, we provide additional insights about the data and the way they are being preprocessed.
@@
 (c) RNA chunks after size filtering
-Figure A.1: Distribution of RNAs or RNA chunks sizes
-In order to work at a more biologically relevant scale, we partition the raw RNAs from PDB files into connected components in order to use these connected components as samples in our different machine learning tasks (except RNA-Ligand in which the data are partitioned in a different way, into binding pockets rather than into connected components).
-Figure 1(c) represents the distribution of the number of residues by connected component chunks.
+Figure 5: Distribution of RNAs or RNA chunks sizes
+A single system, such as rRNA, can gather several independent, non-interacting RNA chains in a single PDB file.
+In order to work at a more biologically relevant scale, we partition the raw RNAs from PDB files into connected components.
+Figure 5(c) represents the distribution of the number of residues by connected components.
 We remove the RNA components of insufficient size for structure-based machine learning, as well as very large components which negatively impact the computational performance of the data loading as well as model training.
-To this end, we filter RNA chunks and only keep the ones having between 15 and 300 nucleotides.
+To this end, we filter RNA chunks and only keep the ones having between 15 and 500 nucleotides.
 Figure 5(c) displays the distribution of RNA chunks sizes after filtering.
 The redundancy removal process we apply is the following.
-First, we perform sequence-based clustering based on a similarity threshold above which RNA fragments are clustered together (relying on sequence similarity metric CD-HIT (Fu et al., 2012)).
+First, we perform sequence-based clustering based on a similarity threshold above which RNA fragments are clustered together (relying on a sequence similarity metric computed using CD-HIT [Fu et al., 2012] clustering program).
 Then, within each cluster, we select the sample with the highest resolution.
-Then we perform structure-based clustering and structure-based redundancy removal (relying on structural similarity metric US-align (Zhang et al., 2022a)) following the same procedure.
-Afterwards, when instantiating the splitters, we perform structure-based clustering with a different threshold to define the clusters which will be required to be grouped either in train, val or test set (which we name ”splitting clustering”).
+Then we perform structure-based clustering and structure-based redundancy removal (relying on the TM-Score, a structural similarity metric, computed using US-align platform [Zhang et al., 2022a]) following the same procedure.
+Afterwards, when instantiating the splitters, we perform structure-based clustering with a different threshold to define the clusters which will be required to be grouped either in train, val or test set (which we call "splitting clustering").
 Since we only select one representative sample per cluster, the number of clusters prefigures the number of samples of the final dataset.
@@
 For this reason, we study the impact of both the redundancy removal and the splitting clustering threshold on the number of clusters generated in the case of RNA-CM task.
-Results are reported in figure A.2.
+Results are reported in figure 6.
 Figure 6(a) displays, for 4 different sequence-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold.
 Here, the redundancy removal thresholds 0.90, 0.80 and 0.70 give the same plot since CD-HIT similarity values are strongly concentrated around 0.5 and 1.
-We finally choose 0.90.
+The chosen value is 0.90.
 Figure 6(b) displays, for 4 different structure-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold.
@@
 (b) Using structure-based redundancy removal
-Figure A.2: Number of clusters after similarity clustering based on the splitting clustering threshold
-A.2 Task-specific data preprocessing
-In this section, we provide additional details about the tasks involving a specific data preparation process.
-A.2.1 RNA-Ligand
-In this task, we selected the three most frequent ligands which were neither modified RNA or DNA residues nor modified protein residues, following the ligand definition proposed by Hariboss (Panei et al., 2022).
+Figure 6: Number of clusters after similarity clustering based on the splitting clustering threshold
+A.1.2 Task-specific data preprocessing
+We provide additional details about the tasks involving a specific data preparation process.
+RNA-Ligand
+We selected the three most frequent ligands which were neither modified RNA or DNA residues nor modified protein residues, following the ligand definition proposed by Hariboss [Panei et al., 2022].
 We only retained the binding pockets binding to one of these three ligands in order to ensure the amount of binding pockets binding to each ligand would be significant enough to enable learning in a multi-class classification framework.
-The three ligands retained are paromomycin (called PAR in the PDB nomenclature), gentamycin C1A (LLL) and aminoglycoside TC007 (8UZ).
-Their structures are displayed in Figure A.3.
+The three ligands retained are paromomycin (called PAR in the PDB nomenclature), gentamicin C1A (LLL) and aminoglycoside TC007 (8UZ).
+Their structures are displayed in Figure 7.
 (a) Paromycin (PAR)
@@
 (c) Aminoglycoside TC007 (8UZ)
-Figure A.3: Structures of the ligands selected for RNA-Ligand task
-A.2.2 RNA-GO
-When building RNA-GO, we explore all the GO-terms of RNAs from the PDB, remove the GO-terms which occur more than 1000 times (these are in fact the GO-terms corresponding to ”structural constituent of ribosome”, ”ribosome” and ”tRNA”).
+Figure 7: Structures of the ligands selected for RNA-Ligand task
+RNA-GO
+When building RNA-GO, we explore all the GO-terms of RNAs from the PDB, remove the GO-terms which occur more than 1000 times (these are in fact the GO-terms corresponding to "structural constituent of ribosome", "ribosome" and "tRNA").
 We also remove the GO-terms which are underrepresented (those which occur less than 50 times in our experiments).
@@
 After this preprocessing, 5 GO-terms are remaining: 0000353 (formation of quadruple SL/U4/U5/U6 snRNP), 0005682 (U5 snRNP), 0005686 (U2 snRNP), 0005688 (U6 snRNP) and 0010468 (regulation of gene expression).
-Appendix B Additional results
-Table A.1: Test performance metrics for various RNA-related tasks.
+A.2 Additional Results
+Table 3: Test performance metrics for various RNA-related tasks.
 Task Test F1-Score Test AUC Test Global Balanced Accuracy Test MCC Test Jaccard
@@
 For RNA_IF, the reported value in “Global Balanced Accuracy” is sequence recovery.
-Table A.2: We compare a standard RGCN using the rnaglib’s task module with various published results using the TR60/TE18 split. Note: Binding site definitions may vary slightly between models.
+Table 4: We compare a standard RGCN using the rnaglib’s task module with various published results using the TR60/TE18 split. Note: Binding site definitions may vary slightly between models.
 Methods MCC AUC
-Rsite2 (Zeng & Cui, 2016) 0.010 0.474
-Rsite (Zeng et al., 2015) 0.055 0.496
-RBind (Wang et al., 2018) 0.141 0.540
-RNAsite_seq (Su et al., 2021) 0.160 0.641
-RNAsite_str (Su et al., 2021) 0.185 0.695
-RNAsite (Su et al., 2021) 0.186 0.703
+Rsite2 [Zeng and Cui, 2016] 0.010 0.474
+Rsite [Zeng et al., 2015] 0.055 0.496
+RBind [Wang et al., 2018] 0.141 0.540
+RNAsite_seq [Su et al., 2021] 0.160 0.641
+RNAsite_str [Su et al., 2021] 0.185 0.695
+RNAsite [Su et al., 2021] 0.186 0.703
 rnaglib RNA-Site 0.113 0.606
-Table A.3: Sequence recovery scores for RNA inverse folding models. We use a standard two layer RGCN part of rnaglib’s task module on the dataset and split published by Joshi et al. (2024)
+Table 5: Sequence recovery scores for RNA inverse folding models. We use a standard two layer RGCN part of rnaglib’s task module on the dataset and split published by Joshi et al. [2024]
 .
-Method Sequence Recovery gRNAde (Joshi et al., 2024) 0.568 Rosetta (Leman et al., 2020b) 0.450 RDesign (Tan et al., 2023) 0.430 FARNA (Alam et al., 2017) 0.321 ViennaRNA (Lorenz et al., 2011) 0.269 rnaglib RNA-IF 0.410
+Method Sequence Recovery gRNAde [Joshi et al., 2024] 0.568 Rosetta [Leman et al., 2020b] 0.450 RDesign [Tan et al., 2023] 0.430 FARNA [Alam et al., 2017] 0.321 ViennaRNA [Lorenz et al., 2011] 0.269 rnaglib RNA-IF 0.410
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The abstract and introduction accurately summarises and introduces our proposal of a benchmarking suite for RNA structure-function modelling.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: We discuss limitations arising from limited structural data, RNA structure idiosyncrasies, as well as the limitating our our experiments to the illustration of our tool’s capabilities.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: We do not provide theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Indeed this information is provided. In fact, our tool ensures that these results can be replicated with minimal effort.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: Yes, all code is provided and the data is publicly hosted. Our tool is well documented and demo and tutorial code is provided for all functionalities.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The paper elaborates in detail on those decisions, especially for dataset processing and splitting.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [N/A]
+  34. Justification: The results described are intended as illustrations of the packages capabilities and are not the main contribution of the paper.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [No]
+  39. Justification: The experiments are based only on small, standard models that do not use significant compute that can, with some patience, be run locally.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44.
+Justification: The paper proposes a benchmarking suite for RNA structure-function modeling.
+Neither the research process, nor the tool have potentially harmful consequences going beyond general research in drug discovery.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: Although down stream applications in drug discovery are discussed, no direct societal impacts result from the research described.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: Our data is all already publicly available through the PDB.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [No]
+  59. Justification: All assets used are properly cited and all data used is open access. For reasons of practicality, not every asset used has its license explicitly mentioned.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: The new assets include extensive documentation for all parts of the functionality and are available through a well maintained and described github repo.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: No human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: No human subjects or crowdsourcing.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: LLMS have been used to assist in writing and development, but their usage is neither important, original, or non-standard with respect to the core method.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@
-A Comprehensive Benchmark for RNA 3D Structure-Function Modeling
+A Comprehensive Benchmark for RNA 3D Structure–Function Modeling
 Abstract
 The relationship between RNA structure and function has recently attracted interest within the deep learning community, a trend expected to intensify as nucleic acid structure models advance.
-Despite this momentum, a lack of standardized, accessible benchmarks for applying deep learning to RNA 3D structures hinders progress.
+Despite this momentum, the lack of standardized, accessible benchmarks for applying deep learning to RNA 3D structures hinders progress.
 To this end, we introduce a collection of seven benchmarking datasets specifically designed to support RNA structure–function prediction.
@@
 To demonstrate the utility of our benchmarks, we report baseline results for all tasks using a relational graph neural network.
-Data and code are available at github.com/cgoliver/rnaglib.
+All code and data, along with documentation are available at rnaglib.org
 1 Introduction
-Recent years have witnessed the advent of deep learning methods for structural biology culminating in the award of the Nobel Prize in Chemistry in 2024.
+Recent years have witnessed the advent of deep learning methods for structural biology, culminating in the award of the Nobel Prize in Chemistry in 2024.
 AlphaFold (Jumper et al., 2021) revolutionized protein structure prediction, equipping the field with millions of new structures.
-Breakthroughs go beyond structure prediction, notably in protein design (Watson et al., 2023; Dauparas et al., 2022), drug discovery (Schneuing et al., 2024; Corso et al., 2022) or fundamental biology (van Kempen et al., 2022).
+Breakthroughs go beyond structure prediction, notably in protein design (Watson et al., 2023; Dauparas et al., 2022), drug discovery (Schneuing et al., 2024; Corso et al., 2022) and fundamental biology (van Kempen et al., 2022).
 A key factor in these breakthroughs was the development of neural encoders that directly model protein structure (Jing et al., 2020; Zhang et al., 2022b; Gainza et al., 2020; Wang et al., 2022).
-These advances in turn are rooted in solid competitions (CASP, CAPRI), and benchmarks (Townshend et al., 2021a; Kucera et al., 2023; Zhu et al., 2022; Jamasb et al., 2024; Notin et al., 2023).
-By setting clear goals, such benchmarks are the foundation for the development of structure encoders and are required to leverage ever increasing structural data availability.
-Yet, to date, structure-function benchmarks have focused on proteins, hampering model development for other biological macromolecules.
+These advances in turn are rooted in robust competitions (CASP, CAPRI) and benchmarks (Townshend et al., 2021a; Kucera et al., 2023; Zhu et al., 2022; Jamasb et al., 2024; Notin et al., 2023).
+By setting clear goals, such benchmarks are the foundation for the development of structure encoders and are required to leverage ever-increasing structural data availability.
+Yet, to date, structure–function benchmarks have focused on proteins, hampering model development for other equally important biological macromolecules.
 Ribonucleic acids (RNAs) are a large family among these macromolecules, supporting biological functions along every branch in the tree of life.
-Besides messenger RNAs which code for proteins, non-coding RNAs are a vast class of molecules which can adopt complex 3D folds (Cech and Steitz, 2014) and take up diverse roles in cellular functions, including gene regulation, RNA processing, and protein synthesis (Statello et al., 2021).
+Besides messenger RNAs which code for proteins, non-coding RNAs are a vast class of molecules which can adopt complex 3D folds (Cech and Steitz, 2014) and play diverse roles in cellular functions, including gene regulation, RNA processing, and protein synthesis (Statello et al., 2021).
 However, our understanding of non-coding RNAs and their functions remains limited.
-This can be largely attributed to the negatively charged nature of RNA backbones, which makes it flexible and limits the availability of high-resolution RNA structures, and imposes significant modeling challenges.
+This can be largely attributed to the negatively charged nature of RNA backbones, which makes them flexible and limits the availability of high-resolution RNA structures, and poses significant modeling challenges.
 Another predominant challenge to a functional understanding of RNA 3D structure lies in the lack of infrastructure for the development and evaluation of function prediction models.
-In this work, we propose a benchmarking suite to act as a facilitating library for the development of structure encoders on RNA.
-Our key contributions include:
+In this work, we propose a benchmarking suite to serve as a facilitating library for the development of structure encoders on RNA.
+Our key contributions are:
   • Seven tasks related to RNA 3D structure that represent various biological challenges.
@@
   • Modular annotators, filters and splitting strategies, both novel and from existing literature, that facilitate the construction of new and custom tasks by other researchers, reducing the time spent on data preparation while ensuring reproducibility of achieved results.
-  • We apply the benchmark to simple architectures to initialize a leaderboard and investigate the effect of splitting and structural context on performance.
+  • We apply the benchmark to simple architectures to initialize a leaderboard and investigate the effect of splitting, structural context, and RNA representation choice on performance.
 2 Related Work
@@
 Classic tasks on proteins appeared independently in unrelated papers, such as GO term and EC number prediction (Gligorijević et al., 2021), fold classification (Hou et al., 2018), binding site detection and classification (Gainza et al., 2020) or binding affinity regression (Wang et al., 2005).
-To our knowledge, ATOM3D (Townshend et al., 2021a) was the first systematic benchmark for molecular systems, albeit heavily focused on proteins.
-Similar, more comprehensive tools were then proposed, such as ProteinShake (Kucera et al., 2023) ProteinWorkshop (Jamasb et al., 2024) and TorchDrug (Zhu et al., 2022), that unify the above tasks and lower the barrier to develop protein structure encoders.
+ATOM3D (Townshend et al., 2021a) was the first systematic benchmark for molecular systems, albeit heavily focused on proteins.
+More comprehensive tools were proposed, such as ProteinShake (Kucera et al., 2023) ProteinWorkshop (Jamasb et al., 2024) and TorchDrug (Zhu et al., 2022), that unify the above tasks and lower the barrier to develop protein structure encoders.
 ProteinGym (Notin et al., 2023) addresses the evaluation of protein mutation effects, while Buttenschoen et al.
@@
 (2024) address protein interactions.
-These works contain notable efforts to scale datasets using predicted structures and to propose biologically appropriate splitting strategies.
+These works are notable efforts to scale datasets using predicted structures and to propose biologically appropriate splitting strategies.
 2.2 RNA Structural Datasets and Benchmarking
-In the realm of RNA 3D structure based modeling infrastructure, three papers propose datasets with the objective of facilitating machine learning.
-They either provide cleaned files in various formats through a web interface (RNAsolo (Adamczyk et al., 2022)), join RNA structures with their corresponding sequence alignment (RNANet (Becquey et al., 2021)) or focus on structure prediction (RNA3DB (Szikszai et al., 2024)).
+In RNA 3D structure-based modeling infrastructure, three papers propose machine learning ready datasets.
+They either provide cleaned files in various formats through a web interface (RNAsolo (Adamczyk et al., 2022)), join RNA structures with their sequence alignment (RNANet (Becquey et al., 2021)) or focus on structure prediction (RNA3DB (Szikszai et al., 2024)).
 Recently, multiple benchmarking suites were introduced for models on RNA sequences (BEACON (Ren et al., 2024)), predicted structures of short RNAs (Xu et al., 2024), or for fitness prediction (Arora et al., 2025).
 None of these methods propose benchmark tasks to compare RNA modeling and learning methods on experimental structures.
+Our effort is integrated within a previously published python package, rnaglib (Mallet et al., 2022) that facilitates the use of RNA structural data with 2.5D graphs.
+This work adds all necessary material to process data and datasets, as well as a set of seven tasks (see Supplementary Section A for a more detailed comparison).
 2.3 Structure-Based RNA Models
-While most deep learning models on RNA focus on secondary structure prediction and sequence-level tasks, some structure-based models have been developed.
+Whilst most deep learning models on RNA focus on secondary structure prediction and sequence-level tasks, some structure-based models were developed.
 RBind (Wang et al., 2018) proposed learning on RNA structures using residue graph representations, followed by others integrating sequence features in the learning (Su et al., 2021; Wang et al., 2023).
-RNAmigos and RNAmigos2 (Oliver et al., 2020; Carvajal-Patiño et al., 2025) proposed incorporating non-canonical interactions in the graph, in conjunction with metric learning pretraining.
+RNAmigos (Oliver et al., 2020) proposed incorporating non-canonical interactions in the graph, in conjunction with metric learning pretraining.
 ARES (Townshend et al., 2021b) applied a tensor field neural network (Thomas et al., 2018) to RNA structure, modeling it as an atomic graph.
-Geometric Vector Perceptron (Jing et al., 2021), an equivariant message passing protein encoder, was adapted to RNA in (Joshi et al., 2024; Tan et al., 2023, 2025; Huang et al., 2024; Wong et al., 2024).
+Geometric Vector Perceptron (Jing et al., 2021), an equivariant message passing protein encoder, was adapted to RNA in (Joshi et al., 2024; Tan et al., 2023; 2025; Huang et al., 2024; Wong et al., 2024).
 EquiRNA (Li et al., 2025) proposed a hierarchical architecture.
-Lately, HARMONY (Xu et al., 2025) introduces a model using multiple RNA modalities.
+Recently, HARMONY (Xu et al., 2025) introduces a model using multiple RNA modalities.
 Despite booming interest in deep learning-based modeling RNA 3D structure, we lack the means to systematically compare models and onboard new practitioners.
@@
 We describe the full process of building a task: data collection, quality filtering, cleanup, and splitting.
-We developed these tools to build our proposed tasks.
-They are ready to use and documented for any practitioner wishing to build his own reproducible task with minimal effort.
-Data collection and annotation
-Our data originates from the RCSB-PDB DataBank (Kouranov et al., 2006), where we fetch all 3D structures containing RNA.
-In some tasks, filter down the set of RNA-containing PDB structures with the subset proposed in (Leontis and Zirbel, 2012) and later referred to as bgsu.
-We annotate each system with RNA-level features such as its resolution, and residue-level features such as their coordinates, the presence of interacting compounds, and the amount of protein atoms in the residues’ vicinities.
-Structure partitioning and quality filters
-RNA molecules present in PDB files exhibit a bimodal distribution over the number of residues (see figures 5(a) and 5(b)).
+Tasks are ready to use and documented for any practitioner wishing to build their own reproducible tasks with minimal effort.
+Data collection and annotation Our data originates from the RCSB-PDB DataBank (Kouranov et al., 2006), where we fetch all 3D structures containing RNA.
+In some tasks, we use the non-redundant structures subset proposed in (Leontis and Zirbel, 2012) and later referred to as bgsu.
+We annotate each system with RNA-level features such as its resolution, and residue-level features such as coordinates, presence of interacting compounds or number of protein atoms in the residues’ vicinities.
+Structure partitioning and quality filters RNA molecules present in PDB files exhibit a bimodal distribution over the number of residues (see Supplementary Figures 5(a) and 5(b)).
 Many systems have less than 300 residues, while some (mostly rRNAs) have several thousands.
-We consider independent RNA fragments from the same PDB file should be treated as multiple systems ; we split each of our systems into connected components in the graph formed by RNA residues and their interactions.
-Concretely, the resulting decrease in system size leads to a reduction in required computing power for downstream steps.
-We implement resolution and size filters, with default cutoff values including systems below 4Å resolution and with a length of 15 to 500 residues.
-In tasks, the upper size cutoff is used to limit computational expense.
-We also provide a protein content filter which removes RNA structures that are heavily structured through an interaction with a protein.
-This crucial filter has been overlooked in most existing RNA structural datasets.
-In addition, we implement the drug-like filter introduced in Hariboss (Panei et al., 2022) for small molecules binding to RNA.
-Redundancy filtering and dataset splitting
-Biological data points are often related in terms of evolutionary history and topological characteristics.
-Moreover, as structures of a given system are often observed in slightly different conditions, it is common to obtain copies with minimal variations, introducing redundancy in the data.
-Insufficient rigor in splitting can thus lead to severe data leaks (Bernett et al., 2024; Volkov et al., 2022).
-Efforts like the PINDER database (Kovtun et al., 2024) have set a standard in the field for model generalization assessment.
-Given a set of RNA molecules $\mathcal{R}$, a similarity metric ${s}$ and a threshold $\theta$, denote as $\mathcal{G}$ the graph $(\mathcal{R},\mathcal{E}^{\theta})$, where $\mathcal{E}^{\theta}:=\{(r_{1},r_{2})\in\mathcal{R}^{2},s(r_{1},r_{2})>\theta\}$.
-We propose a clustering algorithm that returns connected components of $\mathcal{G}$ as clusters.
-This ensures that any pair of points in different clusters has a maximal similarity of $\theta$.
-In our work, we compute sequence-based with CD-HIT (Fu et al., 2012) or structure-based similarity with US-align (Zhang et al., 2022a).
-We propose a redundancy removal algorithm that selects the highest resolution system for each cluster.
-For most tasks, we first apply this algorithm at a sequence similarity threshold of 0.9, and then at a structure similarity threshold of 0.8.
-rnaglib includes a data splitting algorithm that aggregates clusters together.
-To prevent data leakage, we build clusters with a more conservative structural similarity threshold (0.5 unless mentioned otherwise).
-Then, we cast the aggregation of clusters as an integer programming problem, solved using the linear programming tool PuLP (Mitchell et al., 2011).
-This enforces structural dissimilarity between splits for predetermined split size ratios while achieving optimal label stratification.
-Metrics computation
-Given the type of prediction target (e.g.
-binary classification, multi-label, etc.), each task implements a standardized way to compute evaluation metrics.
+We treat RNA fragments from the same PDB file as multiple systems if they do not interact, decreasing system size and reducing required computing power for downstream steps.
+We implement resolution and size filters, with default cutoffs including systems below 4Å resolution and between 15 to 500 residues; the upper size cutoff limits computational expense.
+A novel protein content filter removes RNA structures heavily dependent on protein interactions, a crucial step overlooked in most existing datasets.
+We also include the drug-like filter from Hariboss (Panei et al., 2022) to screen small molecules binding to RNA.
+Redundancy filtering and dataset splitting Rigorous data splitting is often key for structural biology benchmarks such as the PINDER database (Kovtun et al., 2024) for proteins, which has set a standard in the field for model generalization assessment.
+Given a set of RNA molecules $\mathcal{R}$, a similarity metric s{s} and a threshold $\theta$, let $\mathcal{G}$ be the graph connecting similar RNAs, $\mathcal{G}=(\mathcal{R},\mathcal{E}^{\theta})$, where $\mathcal{E}^{\theta}:=\{(r_{1},r_{2})\in\mathcal{R}^{2},s(r_{1},r_{2})>\theta\}$.
+We propose a clustering algorithm that returns connected components of $\mathcal{G}$ as clusters, ensuring that points in different clusters have a maximal similarity of $\theta$.
+We compute similarity s{s} with CD-HIT (Fu et al., 2012) for sequence and with US-align (Zhang et al., 2022a) for structure.
+Building on this, we propose a redundancy removal step by selecting the highest-resolution RNA in each cluster.
+For most tasks, we use a 0.9 sequence similarity threshold and then 0.8 for structural similarity.
+We also propose a data splitting algorithm that aggregates clusters.
+To prevent data leakage, we build dissimilar clusters with a conservative structural similarity threshold (0.5 unless mentioned otherwise).
+Then, we cast the cluster aggregation into splits of given sizes as an integer programming problem, solved with the linear programming tool PuLP (Mitchell et al., 2011); this enforces structural dissimilarity and allows using label stratification.
+Metrics computation Given the type of prediction target (e.g., binary classification, multilabel, etc.), each task implements a standardized way to compute evaluation metrics.
 Each task is evaluated by a recommended metric, specified in Table 1, and a range of additional standard metrics.
-Anatomy of a task
-A “task” is now simply the collection of the aforementioned components with the choices relevant to a particular biological problem.
+Anatomy of a task A “task” is now simply the collection of the aforementioned components with the choices relevant to a particular biological problem.
 That is, each task consists of (i) a dataset drawn from our 3D structure database and processed using our annotations, partitions and filters, (ii) redundancy removing and fixed splits based on sequence, structure or existing literature, and (iii) a well-defined evaluation protocol.
@@
 4 Tasks on RNA 3D Structure
-Our tasks suite explores various dimensions of RNA structural research briefly summarized and illustrated in Figure 1.
+Our tasks suite explores various dimensions of RNA structural research illustrated and briefly summarized in Figure 1.
 We offer seven tasks among which three are based on previous research with published datasets (RNA-IF (Joshi et al., 2024), RNA-Site (Su et al., 2021) and RNA-VS (Carvajal-Patiño et al., 2025)).
 For RNA-IF and RNA-Site, we propose an enhanced, expanded version, resulting in a total of nine separate datasets.
-Armed with an RNA structure, the main focus of each of the subsequent tasks is to either predict aspects of the RNA’s biological function, or to reverse the direction and predict the sequence from which a given structure arises (RNA design).
-We provide an overview of these tasks from a machine learning point of view relating to the type of task and dataset size in Table 1.
+Armed with an RNA structure, the focus of each of the tasks is to predict aspects of the RNA’s biological function, or to reverse the direction and predict the sequence from which a given structure arises (RNA design).
+Further information on task datasets is found in Table 1 and supplement C.
 Task Name Short description
@@
 6. RNA-Ligand Classify an RNA binding site, based on the ligand it binds to.
-7. RNA-VS Predict the affinity of a small molecule for an RNA binding site, and use it for Virtual Screening (VS).
-Figure 1: A graphical overview of the seven tasks included in rnaglib illustrated through a precursor tRNA (4GCW) and a SAM-I riboswitch (5FJC).
-The integrated table contains a short summary of each task.
+7. RNA-VS Predict small molecule-RNA binding site affinity, and use it for Virtual Screening (VS).
+Figure 1: A graphical overview of the seven tasks included in our library, illustrated by a precursor tRNA (4GCW) and a SAM-I riboswitch (5FJC). The integrated table contains short task summaries.
 4.1 RNA-GO: Function Tagging
-(Definition): This is an RNA-level, multi-label classification task where an RNA is mapped to five possible labels representing molecular functions.
-(Context): Effective function prediction models have the capacity to uncover new structure-function connections.
-The Gene Ontology (GO) (Ashburner et al., 2000) was developed to associate a function to a gene, and thus, indirectly to the RNA or protein it encodes.
-It resulted in discrete functional categories called GO terms.
-Predicting the GO term from a protein structure was proposed as a task in (Gligorijević et al., 2021) and has since been regularly used as a benchmark task.
-A manual annotation of RNA sequences with GO terms is available in Rfam (Griffiths-Jones et al., 2003; Ontiveros-Palacios et al., 2025), a database of non-coding RNA families.
-(Processing): We extracted all GO terms corresponding to an RNA structure.
-Then, we filtered these terms based on their frequency, removing the two most frequent labels (ribosomal systems and tRNAs), as well as infrequent ones (less than 50 occurrences).
-We then grouped together GO-terms that were fully correlated, which resulted in five classes.
-Finally, we discarded RNA fragments with less than 15 residues, resulting in 501 systems.
-We split these systems based on a sequence similarity cutoff, following Gligorijević et al.
+(Definition): This is an RNA-level, multilabel classification task where an RNA is mapped to five possible labels representing molecular functions.
+(Context): Function-prediction models have the capacity to uncover new structure-function connections.
+The Gene Ontology (GO) (Ashburner et al., 2000) was developed to associate a function with a gene, and thereby to the RNA or protein it encodes.
+It resulted in functional categories called GO terms, whose prediction from a protein structure was proposed as a task in (Gligorijević et al., 2021), which is now regularly used.
+A manual annotation of RNAs with GO terms is available in Rfam (Griffiths-Jones et al., 2003; Ontiveros-Palacios et al., 2025), a database of non-coding RNA families.
+(Processing): GO terms corresponding to an RNA structure are extracted, then filtered based on their frequency, removing the two most frequent labels (ribosomal systems and tRNAs), and infrequent ones (fewer than 50 occurrences).
+We group together GO-terms that are fully correlated, resulting in five classes.
+Finally, we discard RNA fragments with fewer than 15 residues, resulting in 499 systems.
+We split the data along sequence similarity, following Gligorijević et al.
 (2021).
@@
 (Definition): This is a residue-level, classification task. Given an RNA structure with masked sequence information (keeping only the coordinates of the backbone), we aim to recover the sequence.
-(Context): One avenue for designing new macromolecules is to design a coarse grained tertiary structure first, for instance scaffolding a structural motif, and then design a sequence that will fold into this given structure.
-This second step is denoted as Inverse Folding (IF), since it maps a known structure to an unknown sequence.
+(Context): One avenue for designing new macromolecules is to fix a coarse-grained tertiary structure first, for instance scaffolding a structural motif, and then predict a sequence that will fold into this given structure.
+This second step is called inverse folding (IF), since it maps a known structure to an unknown sequence.
 Inverse folding is a well-established task in protein research with many recent breakthroughs (Watson et al., 2023), with adaptations for RNA with classical models (Leman et al., 2020a).
-Learning based approaches were pioneered by RDesign (Tan et al., 2023), gRNAde (Joshi et al., 2024) and other papers specifically address backbone generation (Anand et al., 2024) and protein-binding RNA design (Nori and Jin, 2024).
-(Processing): We gather all connected components in our 3D structure database having sizes between 15 and 300.
-Then, we remove identical sequences using a redundancy removal step with CD-HIT with a cutoff of 0.99.
-We cluster the data on structural similarity with a threshold of 0.9, and split these while ensuring a maximal similarity of 0.5.
+Learning-based approaches were pioneered by RDesign (Tan et al., 2023), gRNAde (Joshi et al., 2024), while other papers specifically address backbone generation (Anand et al., 2024) and protein-binding RNA design (Nori and Jin, 2024).
+(Processing): We gather all connected components in our database between 15 and 300 nucleotides, then remove identical sequences using CD-HIT redundancy removal with a threshold of 0.9 and split on structural similarity at a threshold of 0.5.
 We also provide datasets and splits from gRNAde (Joshi et al., 2024).
-Our dataset differs by allowing multi-chain systems, a stricter size cutoff, and a duplicates filter, leading to a reduced dataset size.
+Our dataset differs by allowing multichain systems, a stricter size cutoff, and a duplicates filter, leading to a reduced dataset size.
 4.3 RNA-CM: Dynamic Function Modulation
 (Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict which, if any, of its residues are chemically modified.
-(Context): In addition to the four canonical nucleotides, more than 170 modified nucleotides were found to be integrated in RNA polymers (Boccaletto et al., 2022), affecting the functions of a diverse range of ncRNA (Roundtree et al., 2017).
+(Context): In addition to the four canonical nucleotides, more than 170 modified nucleotides have been found in RNA polymers (Boccaletto et al., 2022), affecting the functions of a diverse range of ncRNAs (Roundtree et al., 2017).
 We propose to detect such chemical modifications from subtle perturbations of the canonical geometry of the RNA backbones.
-(Processing): To build this task, we start from the entire dataset, partition it into connected components and apply our size filter.
+(Processing): We start with the entire dataset, partition it into connected components and apply our size filter.
 Then, we filter for systems that include modified residues, relying on PDB annotations.
-We apply our default redundancy removal and splitting strategies, which results in 185 data points.
-Table 1: Description of the ML characteristics of each of our tasks.
-The number in parenthesis refers to the number of classes.
-We include published reference datasets available through rnaglib below their corresponding tasks.
-Task name Level Type Dataset split sizes
-1. RNA-GO RNA Multi-class (5) 349-75-75
-2. RNA-IF Residue Multi-class (4) 1700-448-581
-   gRNAde (Joshi et al., 2024) 11183-528-235
-3. RNA-CM Residue Binary 138-29-30
-4. RNA-Prot Residue Binary 881-189-189
-5. RNA-Site Residue Binary 157-34-33
-   RNASite (Su et al., 2021) 53-6-17
-6. RNA-Ligand Pocket Multi-class (3) 203-43-44
-7. RNA-VS (Carvajal-Patiño et al., 2025) Pocket Regression 304-34-65
+We apply our default redundancy removal and splitting strategies, which results in 197 data points.
 4.4 RNA-Prot: Biological Complex Modeling
-(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether a protein residue is closer than 8Å to any of its residues.
+(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether each residue is within 8Å of a protein residue.
 (Context): RNAs and proteins often bind to form a functional complex.
-Such complexes are involved in crucial cell processes, such as post-transcriptional control of RNAs (Glisovic et al., 2008).
+Such complexes are involved in crucial cellular processes, such as post-transcriptional control of RNAs (Glisovic et al., 2008).
 RNA structure is often, but not always, heavily disrupted upon interaction with a protein.
 (Processing): To build this task, we start from the bgsu non-redundant dataset, partition it into connected components and apply our size and resolution filters.
-We remove systems originating from ribosomes that are fully encapsulated in a complex.
+We remove systems originating from ribosomes that are fully encapsulated within a complex.
 Then, we retain only RNA interacting with a protein and apply our default redundancy removal and splitting strategies, yielding 1251 data points.
 Tasks for RNA Drug Design
-RNA is increasingly recognized as a promising family of targets for the development of novel small molecule therapeutics  (Falese et al., 2021; Haga and Phinney, 2023; Disney, 2019; Abulwerdi et al., 2019).
-Being able to target RNA would drastically increase the size of the druggable space, and propose an alternative for overused protein targets in scenarios where they are insufficient.
-In addition, they represent a therapeutic avenue in pathologies where protein targets are absent, such as in triple-negative breast cancer (Xu et al., 2020).
-The following tasks aim to address questions in RNA small-molecule drug-design.
+RNA is increasingly recognized as a promising class of targets for the development of novel small molecule therapeutics  (Falese et al., 2021; Haga and Phinney, 2023; Disney, 2019; Abulwerdi et al., 2019).
+The ability to target RNA would drastically increase the druggable space, and propose an alternative for overused protein targets in scenarios where they are insufficient.
+RNAs represent a therapeutic avenue in pathologies where protein targets are absent, e.g.
+triple-negative breast cancer (Xu et al., 2020).
+The following tasks address questions in RNA small-molecule drug-design.
 4.5 RNA-Site: Drug Target Detection
-(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether a ligand is closer than 6Å to any of its residues.
-(Context): The classical flow of a structure-based drug discovery pipeline starts with the identification of relevant binding sites, which are parts of the structure likely to interact with ligands, or of particular interest for a phenotypical effect.
-The structure of the binding site can then be used to condition the quest for small molecule binders, for instance using molecular docking (Ruiz-Carmona et al., 2014).
-The framing of this problem as a machine learning task for RNA was introduced in Rbind (Wang et al., 2018).
-(Processing): We provide datasets and predefined splits from RNASite (Su et al., 2021), which are also utilized by other tools in the field such as RLBind (Wang et al., 2023).
-This dataset contains 76 systems obtained after applying stringent clustering on an older version of the PDB.
-In addition to providing the existing dataset existing in the literature, we propose a larger, up-to-date dataset.
-We start by following similar steps as for RNA-Prot (without removal of ribosomal systems).
-We then include a drug-like filter on the small-molecules.
-Finally, we include a protein-content filter, removing systems with more than 10 protein atom neighbors, ensuring that the binding is modulated by RNA only.
-We use the default redundancy removal and splitting, resulting in 458 systems.
+(Definition): This is a residue-level, binary classification task where given an RNA structure, we aim to predict whether each residue is closer than 6Å to a ligand.
+(Context): The classic structure-based drug-discovery pipeline starts with the identification of relevant binding sites, which are parts of the structure likely to interact with ligands, or of particular interest for a phenotypic effect.
+The binding site structure can then be used to condition the search for small molecule binders, for instance using molecular docking (Ruiz-Carmona et al., 2014).
+The framing of this problem as a machine-learning task for RNA was introduced in Rbind (Wang et al., 2018).
+(Processing): We provide datasets and splits from RNASite (Su et al., 2021), which are also utilized by other tools such as RLBind (Wang et al., 2023).
+This dataset contains 76 systems obtained after applying stringent clustering on an earlier version of the PDB.
+In addition to the published dataset, we propose a larger, up-to-date dataset.
+We start by following similar steps as for RNA-Prot (without removal of ribosomal systems), then apply a drug-like filter on the small-molecules.
+Finally, we remove systems with more than 10 protein atom neighbors, ensuring that RNA modulated binding.
+The default redundancy removal and splitting results in 224 systems.
 4.6 RNA-Ligand: Pocket Categorization
-(Definition): This is a binding site-level, multi-class classification task where the structure of an RNA binding site is classified according to the partner it accommodates.
+(Definition): This is a binding site-level, multi-class classification task where the structure of an RNA binding site is classified according to the ligand it binds.
 (Context): Equipped with a binding site, one wants to use its structure to characterize its potential binders.
-On proteins, the Masif-Ligand tasks (Gainza et al., 2020) gathers all binding sites bound to the seven most frequent co-factors, and aims to classify them based on their partner.
-Inspired by this work we propose the RNA-Ligand task.
-To keep sufficiently many examples per task, we only retained the three most frequent classes.
+On proteins, the Masif-Ligand tasks (Gainza et al., 2020) gathers all binding sites bound to the seven most frequent cofactors, and aims to classify them based on their ligands.
+Inspired by this work, we propose the RNA-Ligand task.
+To retain sufficiently many examples per task, we only retained the three most frequent classes.
 This task is less ambitious than training a molecular docking surrogate and can help understanding the potential modulators of a given binding site.
-(Processing): Starting with similar steps as the RNA-Site, we obtain a set of structures that display RNAs in interaction with one or more drug-like small molecules.
-We then proceed to extract the context of the binding pocket by seeding two rounds of breadth-first search with all residues closer than 8Å to an atom of the binder.
-This results in all binding pockets in our database and we group these pockets with a sequence clustering.
-Then, our aim is to find the most frequent ligands binding in non-redundant pockets.
-To that end, we gather a dataset of binding site clusters that bind to one ligand only.
-We then compute the most frequent ligands among this dataset and retain only the top three.
-Subsequently, we discard all binding events to other small molecules, and add clusters that bind only one of the top three ligands to our dataset and split based on structural similarity.
+(Processing): Starting with steps similar to RNA-Site, we obtain a set of structures that display RNAs in interaction with one or more drug-like small molecules.
+We then extract the context of the binding pocket by seeding two rounds of breadth-first search with all residues closer than 8Å to an atom of the binder.
+This results in binding pockets, which we group by sequence clustering.
+To find the most frequent ligands binding in non-redundant pockets, we discard binding sites that bind to several ligands and retain only binding events to the three most frequent ligands in the remaining set.
+We split this final set based on structural similarity.
 4.7 RNA-VS: Drug Screening
-(Definition): This is a binding site-level, regression task.
-Given the structure of an RNA binding site and a small molecule represented as a molecular graph, the goal is to predict the affinity of their binding.
-(Context): Beyond classification into fixed categories, virtual screening aims to score compounds based on their affinity to a binding pocket.
-This task is ubiquitous in drug design as it helps selecting the most promising compounds to be further assayed in the wet-lab.
-Our last task implements a virtual screening task introduced in (Carvajal-Patiño et al., 2025).
-Their model is trained to approximate normalized molecular docking scores.
-Trained models can then be used to rank compounds by their binding likelihood to target sites, hence achieving virtual screening.
+(Definition): This is a binding site-level regression task.
+Given the structure of an RNA binding site and a small molecule represented as a molecular graph, the goal is to predict their binding affinity.
+(Context): Beyond fixed-category classification, virtual screening aims to score compounds based on their affinity to a binding pocket.
+This task is ubiquitous in drug design as it helps select promising compounds to be further assayed in the wet-lab.
+Here, we implement the virtual screening task introduced in (Carvajal-Patiño et al., 2025).
+Their model is trained to approximate normalized molecular docking scores, and can then rank compounds by their binding likelihood to target sites.
 (Processing): The dataset is reproduced from the RNAmigos2 (Carvajal-Patiño et al., 2025) paper.
-The authors curated a list of binding sites in a similar fashion to the one described in the RNA-Site task and clustered the sites using RMScore (Zheng et al., 2019).
+The authors curated a list of binding sites analogously to the RNA-Site task and clustered the sites using RMScore (Zheng et al., 2019).
 All binders found for each cluster are retained as positive examples, and a set of drug-like chemical decoys are added as negative partners.
@@
 Next, we briefly showcase the use of our framework for a simple programmatic access to proposed tasks.
-In Figure 2, we show how practitioners can access our datasets from Python code, automatically downloading them from Zenodo, choosing a representation (in this example a Pytorch Geometric graph) and directly accessing the data in a simple and reproducible fashion.
-Additionally, by passing a single flag (recompute=True) to the task fetcher, users can choose to execute all processing logic from scratch (end-to-end reproducibility).
+In Figure 2, we show how practitioners can access our datasets from Python code, automatically downloading them from Zenodo, choose a representation (in this example a Pytorch Geometric graph) and directly use the data in a simple and reproducible fashion.
+Additionally, by passing a single flag (recompute=True) to the task fetcher, users can choose to execute all processing logic from scratch, ensuring end-to-end reproducibility.
 1from rnaglib.tasks import get_task
@@
 Thanks to the modularity of our tool, additional tasks on RNA can be easily integrated in our framework for future releases.
-This is illustrated in the documentation at rnaglib.org.
+This is illustrated in the documentation at rnaglib.org
 6 Experiments and Results
-We hereby explicitly state that we limit the scope of this work to the creation of a useful deep learning library for RNA 3D structure based modeling.
-We refrain from suggesting novel model architectures; models discussed in this work uniquely serve the goal of illustrating the practical applicability of our library.
-However, we provide a baseline performance for each of our proposed tasks as a reference for future works and conduct some investigations highlighting rnaglib’s use in understanding and building RNA structure-function models.
-All reported results are conducted over three seeds and code is available from github.com/BorgwardtLab/rnaglib-experiments.
+We explicitly limit the scope of this work to the creation of a useful deep-learning library for RNA 3D structure based modeling.
+We refrain from suggesting novel model architectures and instead aim to illustrate the practical applicability of our library.
+Still, we provide a baseline performance for each of our proposed tasks as a reference for future works and demonstrate our library’s use in understanding and building RNA structure-function models.
+All reported experiments were conducted over three seeds, to report error bars in our plots.
+Implementation details are provided in Section F.
+The code to reproduce our model training and our plots is available from github.com/BorgwardtLab/rnaglib-experiments.
+6.1 Data Analysis and Timing
+We characterize our tasks by analyzing the distribution of RNA families within them, primarily using Rfam (Ontiveros-Palacios et al., 2025), the RNA equivalent of Pfam (Mistry et al., 2021).
+However, due to a significant number of unannotated samples in Rfam, we also use the Nucleic Acid Knowledgebase (NAKB) (Lawson et al., 2023), the intended successor to the Nucleic Acid Database (NDB) (Coimbatore Narayanan et al., 2014), which provides broader coverage and additional nucleic acid–specific annotations.
+Across all tasks, we find a high number of tRNAs and rRNAs, which mirrors the distribution of RNA structures in the Protein Data Bank (PDB).
+A complete breakdown of these family and size distributions is available in Appendix C.
+Despite incomplete annotations, our datasets showcase the variety of functional roles played by RNAs, and thereby illustrate the suitability of our tasks for training models generalizable across families and molecular sizes.
+Training a model on a task can be computationally demanding and thus a potential bottleneck.
+Given the modest amount of RNA structural data, we show that training on our tasks is feasible with limited compute, making the suite accessible to a wide variety of users.
+Concrete training times and memory requirements for each task are reported in Appendix D.
+6.2 Redundancy and Splitting
+Biological data points are inherently correlated due to evolutionary history, topological similarities, and minor experimental variations that introduce redundancy.
+Consequently, this data is not independent and identically distributed (i.i.d.).
+This non-i.i.d.
+nature means that commonly used random data splits can be insufficient (Volkov et al., 2022), which has motivated the development of more advanced splitting strategies (Bernett et al., 2024).
+The choice of splitting strategy depends on the desired type of generalization, from simple interpolation to the more challenging extrapolation across dissimilar proteins or RNAs.
+The latter is driven by the ambition to use machine-learning to approximate the universal physical laws that are believed to govern structure-function relationships (Nori et al., 2025).
+To this end, similarity-aware evaluation allows researchers to precisely tune the desired degree of generalization, bridging the gap between random and strict similarity-based splits (Zhang et al., 2025).
+To showcase our library’s utility in choosing suitable splitting approaches, we compare its three main splitting approaches in Figure 3.
+As expected, random splitting often leads to inflated performance metrics.
+However, this effect was less pronounced on our datasets, which we attribute to our initial redundancy removal step mitigating data leakage.
+Indeed, forgoing the redundancy removal for three example tasks markedly increases the performance gap between similarity-based and random splitting approaches.
+Figure 3: Redundancy removal has a dampening effect on data leakage across splitting strategies.
+For each splitting strategy, our library permits continuous tuning of the similarity thresholds, depending on use case and data availability.
+The trade-off between leakage and learning signal can thereby be tweaked by practitioners.
+In Supplementary Figure 7, we plot the performance of our baseline model on the redundant version of RNA-Site, varying the similarity threshold used.
+Again, a more stringent splitting leads to a decreased performance.
+6.3 Model Benchmarking
+We provide two case studies on selected tasks elucidating how our library can assist in choosing modeling modalities.
+We investigate the impact of increasing the depth of our model, as well as the effect of using different RNA representations.
+We focus on RNA-Site, RNA-CM, and RNA-Prot.
+We evaluated three classes of RNA representations reflecting its multi-level structure. 1D sequence models, such as LSTM or Transformers, use only the raw nucleotide sequence.
+2D coarse-grained graphs represent nucleotides as nodes and encode structural information as edges.
+These range from simple base-pair and backbone connections considered as one (2D) or several edge types (2D+) to the 2.5D representation, which adds non-canonical interactions using Leontis-Westhof nomenclature (Leontis and Westhof, 2001).
+Those graphs are processed by a Relational Graph Convolutional Network (RGCN).
+Finally, 3D geometric graphs explicitly model tertiary structure using residue coordinates, along with $k$-nearest neighbors edges (GVP) or with the 2.5D representation edges (GVP-2.5D).
+These graphs are processed by an equivariant graph neural network, the Geometric Vector Perceptron (GVP) Jing et al.
+(2020).
+Further details about the representations and models are provided in Appendix E.
+[…]
+(a) Figure 4a: Performance of different RNA representations: sequence (green), coarse-grained (blue) or full 3D (red).
+(b) Figure 4b: Performance of the coarse grained approach for different RGCN depths.
+Our results clearly indicate that coarse-grained models outperform 3D methods (Figure 4(a)).
+We hypothesize that atomic-level detail is unnecessary for modeling inherently unstable RNA structures, particularly with limited data.
+This is in line with existing findings on predicted structure Xu et al.
+(2024).
+However, our results also show that 3D methods outperform 1D methods, which contrast with the results on predicted structures, where noisy 3D information offered no advantage over sequence data.
+Our use of experimental structures demonstrates the definitive benefit of incorporating 3D structural information over 1D models.
+Moreover, our results demonstrate that base-pairing information is a crucial prior for both 2D and 3D models, consistent with previous results obtained on a virtual screening task on graph representations Oliver et al.
+(2020).
+The 2.5D representation was the top performer, likely due to its combination of a powerful biological prior (non-canonical pairs) with an efficient, sparse graph structure.
+Finally, we investigated the effect of structural context size by varying the depth of our 2.5D network from two to six layers (Figure 4(b)).
+While four layers consistently represented a strong baseline, the optimal depth is task-dependent.
+For instance, predicting protein binding sites (RNA-Prot) appears to require less structural context than other tasks.
 6.4 Baseline Performance
-The baseline performance for each of our proposed tasks is achieved by training Relational Graph Convolutional Networks (RGCN) on graph representations of the tasks’ underlying datasets.
-We choose to use 2.5D graphs, whose nodes represent residues and edges represent both canonical and non-canonical interactions.
-This approach was previously shown to outperform a secondary structure based alternative in a drug-discovery task (Oliver et al., 2020).
-We report our results in Table 2 and additional details on the experimental setup, including hyperparameters, and further metrics can be found in the supplementary Section A.2.
-We observe that on all tasks, our machine learning models are able to capture a meaningful signal that outperforms naive approaches, always predicting the majority class, validating the relevance of our tasks for benchmarking purposes.
-Table 2: Baseline performance reported using our recommended representative metric for each task. We compare to always predicting the majority class.
-Task Metric Majority Score
-RNA-GO Jaccard 0.2 0.396 $\pm$ 0.023
-RNA-IF Sequence Recovery 0.28 0.372 $\pm$ 0.005
-RNA-CM Balanced Accuracy 0.5 0.633 $\pm$ 0.007
-RNA-Prot Balanced Accuracy 0.5 0.556 $\pm$ 0.004
-RNA-Site Balanced Accuracy 0.5 0.655 $\pm$ 0.012
-RNA-Ligand AuROC 0.33 0.650 $\pm$ 0.124
-RNA-VS AuROC 0.5 0.759 $\pm$ 0.005
+Building on those results, we use our best-performing 2.5D approach for each of our proposed tasks to serve as a baseline.
+We report our results in Table 1 and additional details on the experimental setup, including hyperparameters, and further metrics are in the supplementary Section G.
+We also include the state-of-the-art performance for existing tasks.
 Furthermore, we compare our RGCN approach to existing methods on the TR60/TE18 dataset split (Su et al., 2021) for the RNA-Site task and the dataset and splits from Joshi et al.
 (2024) for the RNA-VS task.
-Performance metrics are provided in Supplementary Tables 4 and 5.
-Despite using a significantly simpler model, our approach achieves competitive results, aligning with recent methods while falling short of state-of-the-art performance.
-6.2 Redundancy and Splitting
-To showcase rnaglib’s use in choosing suitable splitting approaches, we compare the three main splitting approaches it provides and present results in Figure 3(a).
-As expected, training a model on randomly split data leads to inflated results in some cases.
-However, the difference is not clear cut 3(a).
-Since our tasks’ datasets already pass a redundancy removal step, we postulate that the reduced redundancy mitigates the effect of leakage.
-Indeed, forgoing the redundancy removal for an example task (RNA-Site) markedly increases the performance difference between similarity-based and random splitting approaches.
-rnaglib permits setting different structure and sequence similarity thresholds for splitting, depending on use case and data availability.
-In Figure 3(b), we plot the performance of our baseline model on the redundant version of RNA-Site, varying the similarity threshold used.
-Our results show that more stringent splitting leads to a decreased, more realistic performance.
-The trade-off between leakage and learning signal can thereby be tweaked by practitioners.
-We point out that structure and sequence thresholds are not directly comparable, since their calculation differs.
-Setting a threshold of 1 results in isolated points for our structure similarity metric, resulting in comparable results to random splitting.
-This is in contrast with our sequence-based similarity metric, whose value is exactly one for highly similar structures.
-Using either a non-redundant dataset or rigorous splitting, alleviates the risk of data leakage, using both maximizes the reliability of the results.
-(a) Figure 3a: Redundancy removal has a dampening effect on data leakage across splitting strategies.
-(b) Figure 3b: In a redundant setting, stricter splitting thresholds reduce data leakage, shown for RNA-Site.
-6.3 Model Benchmarking
-We provide two case studies on selected tasks elucidating how rnaglib can assist in choosing modeling modalities.
-We investigate the impact of increasing the depth of our model, as well as the effect of using different representations.
-We focus on RNA-Site, RNA-CM, and RNA-Prot.
-First, we assess the impact of varying the number of message passing layers to understand the effect of varying structural context size on model performance.
-We train models of various depths (number of layers ranging from two to six) on each task of these three tasks, tuning only the learning rate which we adjust for optimal convergence, keeping the remaining hyperparameters as in Supplementary Table 3.
-Results are presented in Figure 4(a).
-While using four layers generally results in a good performance, the best value varies from one task to another.
-For example, predicting protein binding sites (RNA-Prot) might require less structural context than other tasks.
-[…]
-(a) Figure 4a: RGCN depth
-(b) Figure 4b: Structural detail in input RNAs
-Next, we assess various types of graph representations for RNA structures.
-We contrast 2.5D RNA graphs, whose edges include both canonical and non-canonical interactions, rnaglib’s default representation, with 2D RNA graphs (i.e.
-canonical interactions only).
-Within the latter, we consider a 2D setting, where all canonical interactions are encoded as one edge type, and a 2D+ setting, where canonical Watson-Crick base-pairing and backbone covalent interactions are distinguished (the backbone’s directionality results in a total of three edge types).
-We compare these three representations in Figure 4(b) and observe that including non canonical interactions results in a performance increase across all tasks.
-These results are in accordance with those presented in Oliver et al.
-(2020).
+Performance metrics are in Supplementary Tables 8 and 9.
+Table 1: Baseline performance reported using our recommended representative metric for each task.
+We compare to always predicting the majority class.
+Scores with an asterisk refer to existing literature tasks and are reported directly from their respective publications.
+Task Metric Majority Score Dataset split sizes
+RNA-GO Jaccard 0.2 0.533 $\pm$ 0.013 349-75-75
+RNA-IF Sequence Recovery 0.28 0.354 $\pm$ 0.005 1700-448-581
+   gRNAde Sequence Recovery 0.28 0.568* 11183-528-235
+RNA-CM Balanced Accuracy 0.5 0.661 $\pm$ 0.017 138-29-30
+RNA-Prot Balanced Accuracy 0.5 0.553 $\pm$ 0.007 881-189-189
+RNA-Site Balanced Accuracy 0.5 0.607 $\pm$ 0.005 157-34-33
+   RNASite AuROC 0.5 0.703* 53-6-17
+RNA-Ligand AuROC 0.33 0.650 $\pm$ 0.006 203-43-44
+RNA-VS AuROC 0.5 0.759 $\pm$ 0.005 304-34-65
+Despite using a significantly simpler model, our approach achieves competitive results on those tasks, aligning with recent methods while falling short of state-of-the-art performance.
+We observe that on all tasks, our simple machine learning models are able to capture a meaningful signal that widely outperforms a naive approach of predicting the majority class.
+This confirms that our benchmark tasks are both learnable and challenging, providing a robust foundation for future model development.
 7 Discussion
-We have introduced a versatile and modular library designed to advance the application of deep learning to RNA structural analysis.
-By providing a suite of diverse and well-defined datasets and splits, our library enables robust benchmarking and facilitates the development of innovative computational models.
-It ensures reproducibility and promotes standardized evaluation, fostering confidence in computational findings.
-Its flexibility allows researchers to easily build additional tasks specific to their research needs in a wide range of RNA structural studies.
-In the future, the development of additional tasks, such as the assessment of structural models (Townshend et al., 2021c) and a greater focus on RNA embeddings, presents exciting opportunities.
-As the understanding of the dynamic nature of biological macromolecules evolves, so may the preferred representation for RNA.
-Our tool’s built-in 2.5D-graph representations offer extensive adaptability and have been used to develop recent competitive RNA structure-based models (Wang et al., 2024).
-Nevertheless, the library can easily be extended to incorporate new or improved representations as they emerge.
-In this work, we unify various biological challenges where RNA structure is a key player, and provide direct access to high quality data and benchmarking functionality for deep learning model development and computational biology communities alike.
-Furthermore, the choice of integrating the benchmark in the rnaglib toolkit opens access to non-specialist researchers, and provides robust guarantees for quality and reproducibility.
+We have introduced a versatile and modular library designed to advance deep learning for RNA structural analysis.
+By standardizing datasets, splits, and evaluation as benchmarking tasks, the library enables reproducible comparisons and lowers the barrier for developing new computational models.
+Its standardized workflows foster confidence in results, and its modularity makes it straightforward for researchers to extend the suite with additional tasks.
+Our experiments show that RNA tasks are both learnable and challenging, and that representation choice matters: coarse-grained 2.5D graphs provide strong baselines, while 3D encoders capture additional signal beyond sequence-only models.
+These findings offer practical guidance for model development and underscore the importance of well-founded splitting strategies.
+Looking forward, the benchmark can be extended to new directions such as structural model assessment (Townshend et al., 2021c), transferable RNA embeddings, and the integration of dynamic conformational ensembles.
+The built-in 2.5D representations have already enabled competitive RNA models (Wang et al., 2024) and hold further promise, while the modular design ensures that new and improved representations can be seamlessly integrated as they emerge.
+By consolidating diverse RNA structure–function challenges into a single, extensible library, we create an accessible common foundation for future research in the field.
 We hope that this benchmark will facilitate the development of new deep learning tools in the highly promising area of RNA structural biology.
-Appendix A Technical Appendices and Supplementary Material
-A.1 Details about datasets
-In this appendix, we first detail more extensively the datasets and data preprocessing applied (Section A.1).
-Then, we provide additional results of the models tested on our tasks together with details about training and hyperparameters used to perform the experiments (Section A.2).
-A.1.1 Datasets and general preprocessing
+Reproducibility Statement
+All results can be reproduced with the code provided in the experiments repository.
+All plotting scripts are also available, along with precomputed results files.
+We also provide two simple scripts to retrain all models.
+The code to use the benchmark and the proposed tasks is extensively documented.
+A multitude of tutorials and how-tos exist with regards to all parts of the package.
+The tool can be easily installed as a python package and is maintained on Github.
+Additional and complete information on preprocessing is available in Appendix  B.
+Results can be reproduced with limited compute as specified in Appendix  D.
+Funding
+This project was provided with computing and storage resources by GENCI at IDRIS thanks to the grant 2025-AD010315435R1 on the supercomputer Jean Zay’s V100/A100/ H100 partition.
+Appendix
+In this Appendix, we first clearly differentiate the contribution of the current work with the previously published rnaglib in Section A.
+We then provide analysis of the initial data, a justification of the generic data preprocessing, and details of extra-steps taken for specific tasks (Section B).
+In Section C, we provide a data analysis of the resulting datasets obtained for each of our tasks.
+In Section D, we analyze computational requirements of training models on our benchmarks.
+In Section F, we report additional details about the representation and models used in the paper, as well as about the training parameters used for the models presented in the main body of the paper.
+Finally, we provide additional results of the models tested on our tasks (Section G).
+Appendix A Positioning with regards to rnaglib
+The proposed tool is integrated within a pre-existing rnaglib python package [Mallet et al., 2022]. In this Section we expand on the differences between the two.
+The pre-existing package consists of a database of pre-computed 2.5D graphs, lowering the barrier to entry for people interested in using this representation in their applications.
+In addition, it proposed a set of methods to compare 2.5D graphs using a customized graph-edit distance, plotting scripts, and a kernel to compare 2.5D rooted subgraphs.
+Finally, it proposed a machine-learning model compatible with this representation, with metric-learning pretraining on the kernels, as proposed in RNAmigos Oliver et al.
+[2020].
+Our current submission significantly expands on these fundamentals.
+Notably, we introduce all the infrastructure necessary to build a task.
+We introduce feature computers to easily integrate additional features and annotations, such as RNA language model embeddings.
+We create a representation abstraction and collating routing to go beyond the exclusive use of 2.5D representation.
+We also add a dataset object that can be saved or downloaded from Zenodo, as well as routines to remove redundancy and split a dataset.
+All of these abstract objects are designed to be flexible and modular so that implementing a representation or splitting strategy smoothly fits in our pipeline.
+We then implement concrete tools using this infrastructure.
+Examples include GO term annotators, USAlign splitters or support for sequence, 3D point clouds, voxel grids or 3D graphs representations, and are detailed in Section 3.
+Finally, we propose our benchmark tasks taking advantage of those tools (see Section 4), as well as results leveraging this benchmark (see Section 6).
+Appendix B Datasets Processing
+B.1 General preprocessing
 In this Section, we provide additional insights about the data and the way they are being preprocessed.
-As mentioned in section 3, RNA molecules present in the PDB files exhibit a bimodal distribution over the number of residues.
+As mentioned in Section 3, RNA molecules in the PDB files exhibit a bimodal distribution over the number of residues.
 Figure 5(a) displays the distribution of the RNA sizes (defined as the number of residues) of the RNAs from the PDB.
@@
 Figure 5: Distribution of RNAs or RNA chunks sizes
-A single system, such as rRNA, can gather several independent, non-interacting RNA chains in a single PDB file.
+A single system, such as rRNA, can include several independent, non-interacting RNA chains in a single PDB file.
 In order to work at a more biologically relevant scale, we partition the raw RNAs from PDB files into connected components.
-Figure 5(c) represents the distribution of the number of residues by connected components.
-We remove the RNA components of insufficient size for structure-based machine learning, as well as very large components which negatively impact the computational performance of the data loading as well as model training.
-To this end, we filter RNA chunks and only keep the ones having between 15 and 500 nucleotides.
+Figure 5(c) shows the distribution of the number of residues by connected components.
+We remove the RNA components of insufficient size for structure-based machine learning, as well as very large components which negatively impact the computational performance of data loading as well as model training.
+To this end, we filter RNA chunks and only keep those with between 15 and 500 nucleotides.
 Figure 5(c) displays the distribution of RNA chunks sizes after filtering.
-The redundancy removal process we apply is the following.
-First, we perform sequence-based clustering based on a similarity threshold above which RNA fragments are clustered together (relying on a sequence similarity metric computed using CD-HIT [Fu et al., 2012] clustering program).
+Our redundancy-removal process is as follows.
+First, we perform sequence-based clustering based on a similarity threshold above which RNA fragments are clustered together (using the sequence similarity metric computed using CD-HIT [Fu et al., 2012] clustering program).
 Then, within each cluster, we select the sample with the highest resolution.
-Then we perform structure-based clustering and structure-based redundancy removal (relying on the TM-Score, a structural similarity metric, computed using US-align platform [Zhang et al., 2022a]) following the same procedure.
-Afterwards, when instantiating the splitters, we perform structure-based clustering with a different threshold to define the clusters which will be required to be grouped either in train, val or test set (which we call "splitting clustering").
-Since we only select one representative sample per cluster, the number of clusters prefigures the number of samples of the final dataset.
+We then perform structure-based clustering and structure-based redundancy removal (relying on the TM-Score, a structural similarity metric, computed using US-align platform [Zhang et al., 2022a]) following the same procedure.
+When instantiating the splitters, we perform structure-based clustering with a different threshold to define the clusters which will be required to be grouped either in train, val or test set (which we call "splitting clustering").
+Since we only select one representative sample per cluster, the number of clusters determines the number of samples of the final dataset.
 Therefore, a tradeoff is to be considered between having a large amount of data and discarding redundant RNAs.
-For this reason, we study the impact of both the redundancy removal and the splitting clustering threshold on the number of clusters generated in the case of RNA-CM task.
+For this reason, we study the impact of both redundancy removal and splitting clustering threshold on the number of clusters generated in the case of the RNA-CM task.
 Results are reported in Figure 6.
@@
 In all cases, the structure-based redundancy removal is performed after a sequence-based redundancy removal using a 0.90 threshold.
-In all the experiments, the splitting clustering is performed based on US-align similarity (structural similarity).
+In all experiments, the splitting clustering is performed based on US-align structural similarity.
 (a) Using sequence-based redundancy removal
@@
 Figure 6: Number of clusters after similarity clustering based on the splitting clustering threshold
-A.1.2 Task-specific data preprocessing
+Our library permits setting different structure and sequence similarity thresholds for splitting, depending on use case and data availability.
+In Figure 7, we plot the performance of our baseline model on the redundant version of RNA-Site, varying the similarity threshold used.
+Our results show that more stringent splitting leads to a decreased, more realistic performance.
+The trade-off between leakage and learning signal can thereby be tweaked by practitioners.
+We point out that structure and sequence thresholds are not directly comparable, since their calculation differs.
+Setting a threshold of 1 results in isolated points for our structure similarity metric, resulting in comparable results to random splitting.
+This is in contrast with our sequence-based similarity metric, whose value is exactly one for highly similar structures.
+Using either a non-redundant dataset or rigorous splitting, alleviates the risk of data leakage, and using both maximizes the reliability of the results.
+Figure 7: In a redundant setting, stricter splitting thresholds reduce data leakage, shown for RNA-Site.
+B.2 Task-specific Preprocessing
 We provide additional details about the tasks involving a specific data preparation process.
-RNA-Ligand
-We selected the three most frequent ligands which were neither modified RNA or DNA residues nor modified protein residues, following the ligand definition proposed by Hariboss [Panei et al., 2022].
+RNA-Ligand We selected the three most frequent ligands that were neither modified RNA or DNA residues nor modified protein residues, following the ligand definition proposed by Hariboss [Panei et al., 2022].
 We only retained the binding pockets binding to one of these three ligands in order to ensure the amount of binding pockets binding to each ligand would be significant enough to enable learning in a multi-class classification framework.
@@
 Their structures are displayed in Figure 8.
-(a) Paromycin (PAR)
-(b) Gentamycin C1A (LLL)
+(a) Paramomycin (PAR)
+(b) Gentamicin C1a (LLL)
 (c) Aminoglycoside TC007 (8UZ)
-Figure 7: Structures of the ligands selected for RNA-Ligand task
-RNA-GO
-When building RNA-GO, we explore all the GO-terms of RNAs from the PDB, remove the GO-terms which occur more than 1000 times (these are in fact the GO-terms corresponding to "structural constituent of ribosome", "ribosome" and "tRNA").
+Figure 8: Structures of the ligands selected for the RNA-Ligand task
+RNA-GO When building RNA-GO, we explore all the GO-terms of RNAs from the PDB, remove the GO-terms which occur more than 1000 times (these are in fact the GO-terms corresponding to "structural constituent of ribosome", "ribosome" and "tRNA").
 We also remove the GO-terms which are underrepresented (those which occur less than 50 times in our experiments).
@@
 After this preprocessing, 5 GO-terms are remaining: 0000353 (formation of quadruple SL/U4/U5/U6 snRNP), 0005682 (U5 snRNP), 0005686 (U2 snRNP), 0005688 (U6 snRNP) and 0010468 (regulation of gene expression).
-A.2 Additional Results
-Table 3: Test performance metrics for various RNA-related tasks.
+Appendix C Data Analysis
+Figures 910 show the family distributions of our proposed datasets, annotated with both Rfam and NAKB labels.
+We have chosen Rfam clans and NAKB annotations as the right level of granularity, since they provide a good overview of the present functions, without being overly detailed or too general.
+Across tasks we observe substantial functional diversity, with the exception of RNA-Ligand, which was pre-processed to include only RNAs known to bind one of three ligands.
+Overall, NAKB provides slightly broader coverage of annotated RNA structures compared to Rfam.
+Importantly, the Rfam-annotated RNAs are not a strict subset of those annotated by NAKB, so the true combined coverage is somewhat higher.
+To our knowledge, however, there is no standardized mapping between Rfam clans and NAKB annotations.
+A large fraction of the structures correspond to ribosomal complexes, particularly ribosomal RNAs and tRNAs.
+This bias reflects the composition of the PDB, where many RNA crystal structures originate from RNAs crystallized in complexes with proteins.
+Given that the ribosome is a central molecular machine of the cell and has been extensively studied by structural biologists, its overrepresentation in our data is expected.
+Table 2 reports dataset statistics, highlighting the wide range of RNA sizes present in our collection.
+Taken together, the diversity of RNA functions and sizes represented in our datasets ensures that models trained on them are likely to generalize across RNA types.
+Table 2: We provide statistics on the size distribution of RNA samples in our datasets. Note that depending on the task, a size cutoff at either 300 or 500 nucleotides is applied.
+Task Name Total Nodes Dataset Size Min Nodes Max Nodes Mean Nodes Median Nodes RNAs $\geq$200 RNAs $\geq$200 (%) Num Classes Num Features
+rna_cm 11,208 197 16 482 56.89 40.00 4 2.0 2 4
+rna_go 46,557 499 10 299 93.30 93.00 3 0.6 5 4
+rna_if 164,769 2,729 16 298 60.38 44.00 72 2.6 5 1
+rna_ligand 14,632 290 27 78 50.46 49.00 0 0.0 3 4
+rna_prot 74,257 1,274 16 488 58.29 40.50 36 2.8 2 4
+rna_site 14,430 224 16 478 64.42 47.00 9 4.0 2 4
+Refer to captionRefer to caption
+(a) CM
+Refer to captionRefer to caption
+(b) GO
+Refer to captionRefer to caption
+(c) IF
+Refer to captionRefer to caption
+(d) Ligand
+Refer to captionRefer to caption
+(e) Prot
+Refer to captionRefer to caption
+(f) Site
+Figure 9: Distribution of NAKB functional annotations across task datasets. For each dataset, the top plot excludes unannotated RNAs and the bottom includes them.
+Refer to captionRefer to caption
+(a) CM
+Refer to captionRefer to caption
+(b) GO
+Refer to captionRefer to caption
+(c) IF
+Refer to captionRefer to caption
+(d) Ligand
+Refer to captionRefer to caption
+(e) Prot
+Refer to captionRefer to caption
+(f) Site
+Figure 10: Distribution of Rfam clans across task datasets. For each dataset, the top plot excludes unannotated RNAs and the bottom includes them.
+Appendix D Timing
+We ran our proposed models with a batch size of 8 on 4 Intel i9 CPU cores or on a personal GPU (RTX 3500), and report runtimes and memory requirements in Table 3.
+Given the modest dataset sizes, runtimes and memory use were not extensively optimized but are unlikely to pose practical limitations.
+Depending on the task, training a model takes between a few minutes and one hour.
+Training on a GPU results in a twofold speedup.
+Table 3: An overview of the time and memory requirements of training our models for each provided task.
+Task Name Time/RNA (CPU ms) Time/RNA (GPU ms) Peak GPU Memory (MB)
+RNA_CM 22.4 9.69 60.66
+RNA_Go 23.77 11.78 48.84
+RNA_IF 10.11 5.49 55.40
+RNA_Ligand 21.46 8.65 42.00
+RNA_Prot 21.64 9.70 52.04
+RNA_Site 23.42 9.95 50.77
+RNA_VS 33.55 16.59 41.97
+Appendix E Details about representations and models
+This section provides a comprehensive overview of the models and RNA representations used in order to foster reproducibility.
+E.1 Models
+This subsection provides details about the models used.
+Bidirectional Long Short-Term Memory (BiLSTM) processes sequential data in both forward and backward directions using two separate Long short-term memory (LSTM) layers.
+Unlike standard LSTMs that only consider past context, BiLSTM captures both past and future information by concatenating outputs from both directions.
+This bidirectional approach provides richer contextual representations, making it effective for sequence labeling and classification tasks where the complete sequence is available during processing.
+Transformers are attention-based neural networks that process sequential data in parallel rather than sequentially.
+The core mechanism is self-attention, which allows the model to weigh the importance of different positions in the input sequence when computing representations for each token.
+Unlike RNNs, transformers can process all sequence positions simultaneously, enabling faster training and better capture of long-range dependencies.
+The architecture consists of encoder-decoder blocks with multi-head attention, position encodings, and feed-forward layers.
+Graph convolutional network (GCN) [Kipf, 2016] is a graph neural network architecture proposing to extend convolution operations to graph data by performing message passing on the 1-hop neighborhood of each node at each layer.
+The formula of the GCN is the following:
+$h^{l+1}=\sigma(Ah^{l}W+b)$
+where $h^{l}$ denotes the embedding at the end of layer $l$, $\sigma$ denotes an activation function, $A$ the adjacency matrix of the graph, $W$ the weights of layer $l+1$ and $b$ the bias of layer $l+1$
+Relational graph convolutional network (RGCN) [Schlichtkrull et al., 2018] is a refinement of GCN with different edge types, and distinct learnable weight matrices for each edge type.
+It has the following formula:
+$h^{l+1}=\sigma(W_{0}h^{l}+\sum_{r=1}^{R}W_{r}A^{r}h^{l})$
+where $R$ is the number of edge types and $A^{r}$ the adjacency matrix relative to edge type $r$
+Geometric vector perceptron (GVP) [Jing et al., 2020], initially introduced to learn on protein structures, is an equivariant graph neural network designed to process both scalar and vector node and edge features.
+In particular, it enables to process 3D coordinates in the learning process.
+The GVP equations consist in the update equations of both scalar $s^{l}$ and vector $V^{l}$ features at layer $l+1$:
+$s^{l+1}=\sigma(W_{m}[\|W_{h}V^{l}\|_{2}||s^{l}])$
+$V^{l+1}=\sigma^{+}(\|W_{\mu}W_{h}V^{l}\|_{2})\odot W_{\mu}W_{h}V^{l}$
+where $\sigma$ and $\sigma^{+}$ are two activation functions and $W_{h}$ and $W_{\mu}$ are two learnable linear transformations, $\odot$ denotes row-wise multiplication and $[\cdot||\cdot]$ denotes concatenation.
+E.2 Representations
+We here detail the various representation frameworks proposed in Rnaglib as Representation objects and used in our experiments.
+2D representation.
+RNA is represented as a graph.
+Each node corresponds to a nucleotide and each edge encodes either a backbone connection or a canonical base pair.
+Node features are the one-hot encoding of base identity of nucleotides.
+Edges are not featurized.
+2D+ representation.
+RNA is represented as a graph, very similar to the 2D representation.
+The only difference lies in the presence of three distinct edge types: backbone 3’-5’ connections, backbone 5’-3’ connections and base pairing edges.
+2.5D representation.
+RNA is represented as a graph whose nodes correspond to a nucleotide and edges encode either a backbone connection or a base pair, be it canonical or non-canonical.
+This representation is built using structures coming from the Protein Data Bank (PDB).
+Node features encode the base identity of nucleotides and, optionally, comprise an embedding computed from the RNA language model RNA-FM [Chen et al., 2022].
+This graph has distinct edge types.
+Not only do these edge types distinguish between backbone and base pair connections (as in 2D+ representation), they also account for the geometry of base pair interactions according to Leontis and Westhof categorization.
+3D representations.
+RNA is represented as a graph which nodes represent its nucleotides.
+We provide two possibilities for edge definition: either a $k$-nearest neighbor graph (edges are built between a node and its $k$ nearest nodes in the 3D space) or a graph which edges encode either a backbone connection or a canonical base pair (as in 2D and 2D+ representation).
+In our experiments, we report results for $k$-nearest neighbor graphs with $k=16$.
+Node scalar features encode the base identity of nucleotides and, optionally, comprise an embedding computed from the RNA language model RNA-FM.
+Node vector features are unit vectors from nucleotide i-1 to nucleotide i and from nucleotide i to nucleotide i+1) and, optionally, the base pair geometry according to Leontis and Westhof nomenclature (this option is typically used in our GVP-2.5D representation).
+Edge scalar features include radial basis function encoding of the distance between the two nucleotides forming the edge.
+As edge vector features, we use the unit vector from nucleotide i to nucleotide j, where i and j are the nucleotides represented by the nodes forming the edge.
+By default and in our experiment, the 3D coordinates of the nucleotides are assumed to be the 3D coordinates of their phosphate atoms.
+However, the implementation of the representation is very modular and gives the users the possibility to add custom edge or node features, as well as to use different representative atoms to model nucleotide coordinates.
+Appendix F Training details
+We provide all the code needed to reproduce our training in the two attached repositories.
+The code is accompanied by readme documents describing the step-by-step process to reproduce our results.
+In addition, extensive code documentation is available for all parts of the developed tool.
+For a simple overview, the most important hyperparameters of the GCNs and RGCNs used in our experiments are summarized in table  4.
+All our experiments have been performed using a batch size of 8.
+Table 4: Hyperparameters used for different RNA tasks in GCN and RGCN models.
+Model Representation Layers Hidden Dim Epochs Learning Rate Dropout
+RNA_Ligand 2.5D 3 64 20 $1\times 10^{-3}$ 0.5
+RNA_CM 2.5D 3 128 40 $1\times 10^{-3}$ 0.5
+2D 3 128 40 $1\times 10^{-3}$ 0.5
+2D-GCN 2 256 40 $1\times 10^{-3}$ 0.5
+RNA_Site 2.5D 4 256 40 $1\times 10^{-3}$ 0.5
+2D 2 128 40 $1\times 10^{-4}$ 0.5
+2D-GCN 4 256 100 $1\times 10^{-3}$ 0.5
+RNA_Prot 2.5D 4 128 40 $1\times 10^{-3}$ 0.2
+2D 4 64 40 $1\times 10^{-2}$ 0.2
+2D-GCN 5 256 40 $1\times 10^{-2}$ 0.2
+RNA_IF 2.5D 3 128 100 $1\times 10^{-4}$ 0.5
+RNA_VS 2.5D 3 64/32 100 $1\times 10^{-3}$ 0.2
+RNA_GO 2.5D 3 128 20 $1\times 10^{-3}$ 0.5
+We report below the hyperparameters used to train GVP models.
+In the case of GVP models, we don’t tune a unique hidden dimension hyperparameter but four: the number of hidden node scalar features $h_{s}^{N}$, the number of hidden node vector features $h_{V}^{N}$, the number of hidden edge scalar features $h_{s}^{E}$ and the number of hidden edge vector features $h_{V}^{E}$.
+In the table below, we call Node Dim the $(h_{s}^{N},h_{V}^{N})$ couple and Edge Dim the $(h_{s}^{E},h_{V}^{E})$ couple.
+Table 5: Hyperparameters used for different RNA tasks in GVP models.
+Model Representation Layers Node Dim Edge Dim Epochs Learning Rate Dropout
+RNA_CM GVP 3 (128, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
+GVP-2.5D 3 (32, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
+RNA_Prot GVP 4 (32, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
+GVP-2.5D 4 (32, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
+RNA_Site GVP 6 (32, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
+GVP-2.5D 6 (32, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
+We report below the hyperparameters used for the sequence models (BiLSTM and Transformer).
+Table 6: Hyperparameters used for different RNA tasks in Sequence models.
+Model Model Layers Hidden Dim Epochs Learning Rate Dropout
+RNA_CM BiLSTM 2 128 500 $1\times 10^{-3}$ 0.5
+Transformer 2 32 100 $1\times 10^{-5}$ 0.2
+RNA_Prot BiLSTM 2 64 500 $1\times 10^{-3}$ 0.2
+Transformer 2 128 100 $1\times 10^{-5}$ 0.2
+RNA_Site BiLSTM 2 256 500 $1\times 10^{-3}$ 0.5
+Transformer 8 256 100 $1\times 10^{-5}$ 0.2
+For all binary tasks, we apply sample weighting to the loss in order to account for the class imbalance. The weights attributed to positive samples are given by the following formula:
+$\sqrt{\frac{N_{neg}}{N_{pos}}}$
+where $N_{neg}$ is the number of negative samples in the dataset and $N_{pos}$ is the number of positive samples.
+For 2D-GCN representations only, we directly use the negative-to-positive ratio as sample weights instead of its square root because we found this reinforcement of the weighting to ease training in this setting.
+Appendix G Additional Results
+G.1 Additional metrics for the reported models
+Table 7: Test performance metrics for various RNA-related tasks. For RNA_IF, the reported “Global Balanced Accuracy” corresponds to sequence recovery.
 Task Test F1-Score Test AUC Test Global Balanced Accuracy Test MCC Test Jaccard
-RNA_Ligand 0.2771 0.6751 0.4678
-RNA_CM 0.1957 0.7393 0.6615 0.1695
-RNA_Site 0.3346 0.5929 0.6309 0.3098
-RNA_Prot 0.4545 0.6654 0.6254 0.2469
-RNA_IF 0.3326 0.6201 0.3523 0.1319
-RNA_VS 0.855
-RNA_GO 0.4074 0.8406 0.7067 0.3167
-Hyperparameters used:
-RNA_Ligand: n_layers=4, hidden_dim=128, lr=0.00001, dropout=0.5
-RNA_CM: n_layers=3, hidden_dim=128, lr=0.001, dropout=0.5
-RNA_Site: n_layers=4, hidden_dim=256, lr=0.001, dropout=0.5
-RNA_Prot: n_layers=4, hidden_dim=64, lr=0.01, dropout=0.2
-RNA_IF: n_layers=3, hidden_dim=128, lr=0.0001, dropout=0.5
-RNA_VS: n_layers=3, hidden_dim=64/32, lr=0.001, dropout=0.2
-RNA_GO: n_layers=3, hidden_dim=64, lr=0.001, dropout=0.5
-For RNA_IF, the reported value in “Global Balanced Accuracy” is sequence recovery.
-Table 4: We compare a standard RGCN using the rnaglib’s task module with various published results using the TR60/TE18 split. Note: Binding site definitions may vary slightly between models.
+RNA_Ligand 0.2771 0.6751 0.4678 - -
+RNA_CM 0.1957 0.7393 0.6615 0.1695 -
+RNA_Site 0.3346 0.5929 0.6309 0.3098 -
+RNA_Prot 0.4545 0.6654 0.6254 0.2469 -
+RNA_IF 0.3326 0.6201 0.3523 0.1319 -
+RNA_VS - 0.855 - -
+RNA_GO 0.4074 0.8406 0.7067 - 0.3167
+G.2 Results of the reported models on pre-existing benchmarks
+For some of our tasks, versions have already been published in the literature, which is why we include a literature version in addition to our own improved version into our library.
+While the dataset may be different from ours either in size, or in stringency of splitting criteria, we chose to include them into our tool so that an easy comparison of newly developed models to existing work is possible.
+To showcase this, we compare some simple models trained with our library on these literature tasks with existing works.
+While we do not achieve state-of-the-art, our RGCN models are competitive, despite being much simpler and smaller than the published methods.
+This is an indication of the suitability of graph representations for RNA structure-function modeling, although further research on this is needed.
+Table 8: We compare a standard RGCN using the rnaglib task module with various published results using the TR60/TE18 split. Note: Binding site definitions may vary slightly between models.
 Methods MCC AUC
@@
 RNASite [Su et al., 2021] 0.186 0.703
-rnaglib RNA-Site 0.113 0.606
+Our library RNA-Site 0.113 0.607
+We compare our simple model on the literature RNA-Site task to other published models evaluated on the same data in Table  8.
+RNASite [Su et al., 2021] combines sequence and structural information based on topological and graph features of the structure into a random forest and thereby yields the best score.
+Older tools like Rsite2 [Zeng and Cui, 2016] are based on secondary structure only and deliver inferior performance.
+This is a clear indication that incorporating 3D structural information is essential for solid performance on this benchmark.
+Similarly, we also benchmark our simple RGCN architecture on the literature version of the RNA-IF task.
+There, Joshi et al.
+[2024] achieve the highest sequence recovery with their model gRNAde.
+gRNAde is a geometric deep-learning pipeline using an SE(3) equivariant encoder-decoder.
+This approach yields higher results as physics based models like Rosetta or the contrastive learning approach RDesign.
+Remarkably, a simple RGCN performs similarly to the latter and opens doors for further development of graph-based RNA design models.
+We report the sequence recovery scores in Table  9.
 Table 9: Sequence recovery scores for RNA inverse folding models. We use a standard two layer RGCN part of rnaglib’s task module on the dataset and split published by Joshi et al. [2024]
 .
-Method Sequence Recovery gRNAde [Joshi et al., 2024] 0.568 Rosetta [Leman et al., 2020b] 0.450 RDesign [Tan et al., 2023] 0.430 FARNA [Alam et al., 2017] 0.321 ViennaRNA [Lorenz et al., 2011] 0.269 rnaglib RNA-IF 0.410
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The abstract and introduction accurately summarises and introduces our proposal of a benchmarking suite for RNA structure-function modelling.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9.
-Justification: We discuss limitations arising from limited structural data, RNA structure idiosyncrasies, as well as the limitating our our experiments to the illustration of our tool’s capabilities.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: We do not provide theoretical results.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: Indeed this information is provided. In fact, our tool ensures that these results can be replicated with minimal effort.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: Yes, all code is provided and the data is publicly hosted. Our tool is well documented and demo and tutorial code is provided for all functionalities.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: The paper elaborates in detail on those decisions, especially for dataset processing and splitting.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [N/A]
-  34. Justification: The results described are intended as illustrations of the packages capabilities and are not the main contribution of the paper.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [No]
-  39. Justification: The experiments are based only on small, standard models that do not use significant compute that can, with some patience, be run locally.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44.
-Justification: The paper proposes a benchmarking suite for RNA structure-function modeling.
-Neither the research process, nor the tool have potentially harmful consequences going beyond general research in drug discovery.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: Although down stream applications in drug discovery are discussed, no direct societal impacts result from the research described.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: Our data is all already publicly available through the PDB.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [No]
-  59. Justification: All assets used are properly cited and all data used is open access. For reasons of practicality, not every asset used has its license explicitly mentioned.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: The new assets include extensive documentation for all parts of the functionality and are available through a well maintained and described github repo.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: No human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: No human subjects or crowdsourcing.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: LLMS have been used to assist in writing and development, but their usage is neither important, original, or non-standard with respect to the core method.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Method Sequence Recovery gRNAde [Joshi et al., 2024] 0.568 Rosetta [Leman et al., 2020b] 0.450 RDesign [Tan et al., 2023] 0.430 FARNA [Alam et al., 2017] 0.321 ViennaRNA [Lorenz et al., 2011] 0.269 Our library RNA-IF 0.410
+This leads us to conclude that graph representations of RNA offers new opportunities for powerful structure-function models for RNA.
+We hope that our library will provide the necessary foundation for that development.